### v6

# Tree embeddings and nonuniqueness in site percolation

Abstract.
We prove a nonuniqueness theorem for Bernoulli site percolation on properly embedded planar graphs, and we obtain a general connectivity principle beyond planarity. Let $G$ be an infinite connected graph properly embedded in ${\mathbb{R}}^{2}$ with minimum degree at least $7$. Then
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2},$
and for every
$p\in\bigl(p_{c}^{\mathrm{site}}(G),\,1-p_{c}^{\mathrm{site}}(G)\bigr),$
Bernoulli(p)(p) site percolation on $G$ has almost surely infinitely many infinite open clusters. In particular, this verifies a conjecture of Benjamini and Schramm for properly embedded planar graphs.
The core new ingredient is an explicit embedded-tree separation mechanism for planar nonuniqueness. We construct embedded trees and an embedded forest whose separation properties yield exponential decay of two-point connection probabilities in the matching graph. To treat the high-density regime, we introduce a binary-tree version of uniform percolation and prove stability of infinite clusters under edge additions, without any bounded-degree assumption.
Beyond the planar theorem, we prove a general lower bound on two-point connectivity under uniqueness for arbitrary infinite locally finite graphs. As a consequence, if
$p_{c}^{\mathrm{site}}(G)<p<p_{\mathrm{conn}}(G),$
then Bernoulli site percolation on $G$ has almost surely infinitely many infinite open clusters.
1. Introduction
1.1. Overview and Main Results
A central problem in percolation theory is to understand the number of infinite clusters in the supercritical phase. On transitive graphs this question is by now classical: for Bernoulli percolation, Aizenman, Kesten and Newman [2] proved that the number of infinite clusters is almost surely 0, 11, or $\infty$, while Burton and Keane [7] and Gandolfi, Keane and Newman [10] established uniqueness throughout the supercritical phase for large classes of amenable graphs. In contrast, on non-amenable graphs one expects a genuine nonuniqueness phase, and a central conjecture of Benjamini and Schramm [3] predicts that $p_{c}<p_{u}$ for every non-amenable quasi-transitive graph.
Planarity offers a different route to nonuniqueness, through duality and matching-graph ideas. In quasi-transitive planar graphs these methods lead to a rather complete picture; see, for example, [5, 22, 18, 12, 13]. The situation is much less understood without quasi-transitivity, where symmetry-based tools such as mass transport are unavailable. The purpose of this paper is to show that, in this nonsymmetric planar setting, one can nevertheless prove nonuniqueness throughout an entire interval by a purely geometric argument.
We consider i.i.d. Bernoulli site percolation on infinite connected planar graphs admitting proper embeddings in ${\mathbb{R}}^{2}$, that is, embeddings for which every compact subset of ${\mathbb{R}}^{2}$ intersects only finitely many vertices and edges. For such a graph $G$, write $p_{c}^{\mathrm{site}}(G)$ and $p_{u}^{\mathrm{site}}(G)$ for the critical and uniqueness thresholds:
(1.1) $p_{c}^{\mathrm{site}}(G)$ $:=\inf\Bigl\{p\in[0,1]:{\mathbb{P}}_{p}\bigl(\text{$G$ has an infinite open cluster}\bigr)=1\Bigr\},$
(1.2) $p_{u}^{\mathrm{site}}(G)$ $:=\inf\Bigl\{p\in[0,1]:{\mathbb{P}}_{p}\bigl(\text{$G$ has a unique infinite open cluster}\bigr)>0\Bigr\}.$
Clearly $p_{c}^{\mathrm{site}}(G)\leq p_{u}^{\mathrm{site}}(G)$. When the inequality is strict, there is a nonuniqueness phase in which infinitely many infinite clusters are expected to appear.
Benjamini and Schramm [3, Conjecture 7] proposed the following planar degree-$7$ picture.
Conjecture 1.1.
Let $G$ be an infinite connected planar graph with minimum degree at least $7$. Then
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2},$
and for every
$p\in\bigl(p_{c}^{\mathrm{site}}(G),\,1-p_{c}^{\mathrm{site}}(G)\bigr),$
Bernoulli site percolation on $G$ has almost surely infinitely many infinite open clusters.
In quasi-transitive planar graphs, nonuniqueness can be recovered from matching-graph duality and symmetry arguments. Without quasi-transitivity, however, these tools are no longer available. Haslegrave and Panagiotou [16] proved that
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2}$
under the minimum-degree assumption and the proper embedding assumption, but the corresponding nonuniqueness statement remained open in this level of generality.
Our main result establishes the full conjectural picture for properly embedded planar graphs.
Theorem 1.2.
Let $G$ be an infinite connected graph properly embedded in ${\mathbb{R}}^{2}$, and assume that every vertex of $G$ has degree at least $7$. Then
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2}.$
Moreover, for every
$p\in\bigl(p_{c}^{\mathrm{site}}(G),\,1-p_{c}^{\mathrm{site}}(G)\bigr),$
i.i.d. Bernoulli(p)(p) site percolation on $G$ has almost surely infinitely many infinite open clusters.
The point of Theorem 1.2 is conceptual as well as quantitative. It shows that planar nonuniqueness can be proved in the absence of transitivity by an explicit geometric mechanism. In particular, the proof does not rely on quasi-transitivity, mass transport, or bounded-degree assumptions.
1.2. The Geometric Setting
The geometric content of the problem enters through the proper embedding. In the quasi-transitive planar setting, duality and matching-graph arguments can often be used at a global level and then converted, via symmetry, into statements about the number of infinite clusters. In the present setting there is no such symmetry, so the embedding itself must supply the mechanism that replaces it.
Proper embeddings play two roles in our argument. First, they allow us to work with finite faces in a locally finite planar environment and to pass naturally to the matching graph. Informally, the matching graph $G_{*}$ is obtained from $G$ by adding edges between vertices that lie on a common finite face. It is the site-percolation analogue of the planar dual, but here it must be used quantitatively rather than through symmetry identities alone.
Second, proper embeddings make it possible to construct explicit separating objects in the plane. The embedded trees and forests built in this paper are not merely auxiliary subgraphs: they provide the geometric barriers that force long connections to cross many disjoint regions. This separation mechanism is the source of the quantitative decay estimates that later drive the nonuniqueness argument.
A further point is that the graphs considered here need not be quasi-transitive and may have nontrivial end structure. In particular, one cannot reduce the analysis to a homogeneous large-scale geometry. The proof therefore has to extract enough rigidity directly from the embedding and from the degree assumption.
1.3. Proof Strategy
We now describe the architecture of the proof of Theorem 1.2. The argument proceeds in three steps, each addressing a different obstruction to nonuniqueness.
The first step is to build an explicit embedded tree $T\subseteq G$ with controlled local geometry. This tree already captures enough expansion to imply
$p_{c}^{\mathrm{site}}(G)\leq p_{c}^{\mathrm{site}}(T)<\tfrac{1}{2}.$
Moreover, the tree structure yields nonuniqueness near $p=\tfrac{1}{2}$. The point is not only to compare critical parameters, but to produce a rigid geometric object whose branches can later be organized into separating families.
Our embedded-tree construction should be compared with the spanning-forest theorem of Benjamini and Schramm [4], which shows that graphs with positive Cheeger constant contain trees with positive Cheeger constant. The present construction is of a different nature: it is explicit, adapted to the planar embedding, and designed to produce left/right separation in the plane. This additional geometric control is what later yields a disjoint embedded forest and exponential decay of two-point connection probabilities in the matching graph.
The second step is to upgrade this construction to an embedded forest of pairwise disjoint trees. The role of this forest is to separate distant vertices in a quantitative way. Roughly speaking, any long connection in the matching graph must cross many disjoint barriers created by the forest, and this leads to exponential decay of two-point connection probabilities in $G_{*}$. This gives control of the part of the supercritical regime in which long-range connectivity in the matching graph is still sufficiently weak.
The third step treats the complementary high-density regime. At this point one needs to understand how infinite clusters behave under the addition of edges, and this becomes delicate without a bounded-degree assumption. To overcome this difficulty we introduce a binary-tree version of uniform percolation and prove a stability statement for infinite clusters under edge additions. This allows us to transfer nonuniqueness information from the matching graph back to the original graph.
A key probabilistic input enters through a general lower bound on two-point connectivity under uniqueness. Let $\mathcal{A}_{1}$ denote the event that there exists a unique infinite open cluster. We prove the following statement, valid for arbitrary infinite connected locally finite graphs.
Proposition 1.3.
Let $G=(V,E)$ be an infinite, connected, locally finite graph. Then for every $u,v\in V$,
${\mathbb{P}}_{p}(u\leftrightarrow v)\geq{\mathbb{P}}_{p}(u\leftrightarrow\infty)\,{\mathbb{P}}_{p}(v\leftrightarrow\infty)\,{\mathbb{P}}_{p}(\mathcal{A}_{1}).$
This yields a general criterion for nonuniqueness. Define
(1.3) $p_{\mathrm{conn}}(G):=\sup\Bigl\{p\in(0,1):\ \lim_{n\to\infty}\sup_{\begin{subarray}{c}x,y\in V\\ d_{G}(x,y)\geq n\end{subarray}}{\mathbb{P}}_{p}(x\leftrightarrow y)=0\Bigr\}.$
Corollary 1.4.
Let $G=(V,E)$ be an infinite, connected, locally finite graph. If
$p_{c}^{\mathrm{site}}(G)<p<p_{\mathrm{conn}}(G),$
then Bernoulli site percolation on $G$ has almost surely infinitely many infinite open clusters.
The criterion in Corollary 1.4 is completely independent of planarity and transitivity and may be viewed as a general probabilistic principle: uniqueness of the infinite cluster forces a nontrivial lower bound on long-range two-point connectivity. In the planar setting, the embedded-forest construction supplies the required decay estimate in the matching graph, and Corollary 1.4 converts this quantitative input into nonuniqueness. Theorem 1.2 then follows by combining this with the stability mechanism described above.
1.4. Relation to Previous and Concurrent Work
On quasi-transitive planar graphs, nonuniqueness phenomena can often be analyzed through duality and matching-graph identities; see, for example, [5, 22, 18, 12, 13]. The present paper addresses a different regime, where quasi-transitivity is absent and the main issue is to replace symmetry-based tools by explicit geometric constructions.
For properly embedded planar graphs with minimum degree at least $7$, Haslegrave and Panagiotou [16] proved the inequality
(1.4) $p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2}.$
Our contribution is a new proof of (1.4), and to establish the corresponding nonuniqueness statement throughout
$\bigl(p_{c}^{\mathrm{site}}(G),\,1-p_{c}^{\mathrm{site}}(G)\bigr)$
for this class of graphs.
After earlier versions of this work, Glazman, Harel and Zelesko [11] proved a general $0/\infty$ theorem for planar percolation processes, covering Bernoulli site percolation at $p\leq\tfrac{1}{2}$. Subsequent work [20] shows that, in full generality, Conjecture 7 is false, although under a natural countability assumption on end-equivalence classes one recovers the same nonuniqueness interval. Against this background, the main novelty of the present paper is to identify an explicit geometric route to planar nonuniqueness without transitivity, based on embedded trees, planar separation, and stability under edge additions.
1.5. Organization of the Paper
Section 2 collects background material and notation. In Section 3 we construct the embedded tree and prove
$p_{c}^{\mathrm{site}}(G)<\tfrac{1}{2},$
together with nonuniqueness in the interval
$\bigl(p_{c}^{\mathrm{site}}(T),\,1-p_{c}^{\mathrm{site}}(T)\bigr).$
Section 4 develops the general connectivity criterion for arbitrary locally finite graphs, including the lower bound on two-point connectivity under uniqueness and its nonuniqueness consequence. In Section 5 we build the embedded-forest separation scheme and derive exponential decay of two-point connection probabilities in the matching graph; combined with the results of Section 4, this yields nonuniqueness in the intermediate regime. Section 6 reduces the near-critical high-density regime to a polygon-counting hypothesis by analyzing end structure and infinite 0-*-clusters. Section 7 verifies this hypothesis for triangulations. Finally, Section 8 introduces the binary-tree version of uniform percolation and a triangulation/stability argument, completing the proof of Theorem 1.2 for general properly embedded planar graphs.
2. Background and Notation
[…]
  • removing loops; and
  • removing multiple edges while keeping exactly one edge between each unordered pair of distinct vertices.
[…]
A walk (or path) of $G$ is an alternating finite or infinite sequence $(\dots,v_{0},e_{0},v_{1},e_{1},\dots)$ with $e_{i}=\langle v_{i},v_{i+1}\rangle$. Since our graphs are simple, we often refer to a walk by its vertex sequence.
A walk is closed if it can be written as $(v_{0},v_{1},\ldots,v_{n})$ with $v_{0}=v_{n}$. A walk is self-avoiding if it visits no vertex more than once. A cycle (or simple cycle) is a self-avoiding closed walk $C=(v_{0},v_{1},\dots,v_{n},v_{0})$.
Once $G$ is embedded in ${\mathbb{R}}^{2}$, a face is a maximal connected component of ${\mathbb{R}}^{2}\setminus G$. Faces may be bounded or unbounded. A face is finite if it is bounded and its boundary consists of finitely many edges.
Definition 2.1.
Let $G=(V,E)$ be an infinite, connected, locally finite, simple, planar graph and fix an embedding of $G$ into the plane. The matching graph $G_{*}=(V,E_{*})$ has the same vertex set as $G$, and for distinct $u,v\in V$ we declare $\langle u,v\rangle\in E_{*}$ if and only if
  1. (1)
    $\langle u,v\rangle\in E$; or
  2. (2)
    $\langle u,v\rangle\notin E$, but $u$ and $v$ share a finite face in $G$.
Definition 2.2.
Let $G=(V,E)$ be a graph. An end of $G$ is a map $\psi$ assigning to every finite set of vertices $K\subseteq V$ a connected component $\psi(K)$ of $G\setminus K$, such that $\psi(K)\subseteq\psi(K^{\prime})$ whenever $K^{\prime}\subseteq K$.
The number of ends of a connected graph is the supremum, over finite subgraphs $K$, of the number of infinite components of $G\setminus K$.
Assume $G$ is connected and properly embedded in the plane. The boundary of a finite face is a closed walk; its degree $|f|$ is the number of steps of that walk. (A vertex or an edge may be visited multiple times by the boundary walk, so $|f|$ may exceed the number of distinct boundary vertices.)
For $u,w\in V$ we write $u\sim w$ if $u$ and $w$ are adjacent. If $v$ is a vertex and $f$ is a face, we write $v\sim f$ when $v$ lies on the boundary of $f$.
[…]
3. Embedded Trees
The goal of this section is to show that every infinite, connected graph properly embedded into the plane with vertex degree at least 7 possesses a tree as a subgraph in which every vertex other than the root vertex has degree $3$ or $4$. We start with geometric properties of planar graphs.
Definition 3.1.
Let $G=(V,E)$ be a locally finite planar graph. Let $v\in V$ and $e_{1},\ldots,e_{d}$ be all the incident edges of $v$ in cyclic order. For $1\leq i\leq d-1$, let $f_{i}$ be the face shared by $e_{i}$ and $e_{i+1}$. Let $f_{d}$ be the face shared by $e_{d}$ and $e_{1}$. Let $f:f\sim v$ denote all the faces $f_{1},\ldots,f_{d}$; note that one face may appear multiple times in $f_{1},\ldots,f_{d}$; in that case, it also appears multiple times in $f:f\sim v$. Define the curvature $\kappa(v)$ at a vertex $v\in V$ to be
$2\pi-\sum_{f:f\sim v}\frac{|f|-2}{|f|}\pi.$
Lemma 3.2 (Combinatorial Gauss–Bonnet for a cycle).
Let $G$ be a locally finite graph properly embedded in $\mathbb{R}^{2}$. Let $C$ be a simple cycle with $n$ vertices, and let $F_{C}$ be the set of (finite) faces contained in the bounded component of $\mathbb{R}^{2}\setminus C$. Let $V_{C}^{\circ}$ be the set of vertices strictly inside $C$. Then
(3.1) $\sum_{z\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim z}\frac{|f|-2}{|f|}\pi=(n-2)\pi+\sum_{z\in V_{C}^{\circ}}\kappa(z).$
In particular, if $\kappa(z)\leq 0$ for all $z\in V$, then
(3.2) $\sum_{z\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim z}\frac{|f|-2}{|f|}\pi\leq(n-2)\pi.$
Proof.
Let $m:=|F_{C}|$. Let $s$ be the number of vertices strictly inside $C$ and let $t$ be the number of edges strictly inside the bounded region enclosed by $C$. Since each interior edge is incident to two faces in $F_{C}$ and each boundary edge of $C$ is incident to exactly one face in $F_{C}$, we have
$\sum_{f\in F_{C}}|f|=2t+n.$
Hence
$\sum_{f\in F_{C}}(|f|-2)\pi=(2t+n-2m)\pi.$
On the other hand,
$\sum_{f\in F_{C}}(|f|-2)\pi=\sum_{z\in V_{C}^{\circ}}\sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi+\sum_{z\in C\cap V}\sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi.$
For $z\in V_{C}^{\circ}$, all faces incident to $z$ lie in $F_{C}$, hence by Definition 3.1,
$\sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi=2\pi-\kappa(z).$
Therefore
$\sum_{z\in C\cap V}\sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi=(2t+n-2m)\pi-\sum_{z\in V_{C}^{\circ}}(2\pi-\kappa(z)).$
Finally, Euler’s formula for the planar map inside $C$ gives
$(s+n)-(t+n)+m=1\quad\Rightarrow\quad t=s+m-1.$
Substitute this into the previous identity to get (3.1). The inequality (3.2) follows immediately if $\kappa\leq 0$. ∎
Remark 3.3.
Recall that the hyperbolic plane ${\mathbb{H}}^{2}$ has constant curvature 1-1. Let $C$ be a polygon in ${\mathbb{H}}^{2}$ with vertex set $\{v_{1},\ldots,v_{n}\}_{v_{i}\in{\mathbb{H}}^{2}}$, and edges $\{\langle v_{i},v_{i+1}\rangle\}_{1\leq i\leq n}$ ($v_{n+1}:=v_{1}$) such that $\langle v_{i},v_{i+1}\rangle$ is the geodesic in ${\mathbb{H}}^{2}$ joining $v_{i}$ and $v_{i+1}$. Let $R_{C}$ be the bounded region enclosed by $C$. Then by the Gauss-Bonnet formula of ${\mathbb{H}}^{2}$ we have
(3.3) $\sum_{i\in[n]}\mathrm{internal\ angle\ of}\ C\ \mathrm{at}\ v_{i}=(n-2)\pi-\mathrm{Area}(R_{C})\leq(n-2)\pi.$
In general computing the internal angle at each $v_{i}$ in (3.3) is not straightforward. However, (3.1) is expressed in terms of face degrees instead of internal angels; and appears to be easier to verify in various situations. The Gauss-Bonnet formula was used to study random walks on planar graphs; see [31, 30].
Lemma 3.4.
Let $G=(V,E)$ be a properly embedded planar graph and assume $\kappa(v)\leq 0$ for all $v\in V$. Let $(v_{0},v_{1},\dots)$ be a (finite or infinite) walk in $G$. For each index $n$ for which both $v_{n-1}$ and $v_{n+1}$ are defined, set $e_{n}^{-}:=\langle v_{n-1},v_{n}\rangle$ and $e_{n}^{+}:=\langle v_{n},v_{n+1}\rangle$.
Let $F_{L}(n)$ (resp. $F_{R}(n)$) be the multiset of faces incident to $v_{n}$ encountered when turning from $e_{n}^{-}$ to $e_{n}^{+}$ around $v_{n}$ counterclockwise (resp. clockwise), where faces are counted with multiplicity as in Definition 3.1. Assume that for every such $n$,
(3.4) $\min\Big\{\sum_{f\in F_{L}(n)}\frac{|f|-2}{|f|}\pi\;,\;\sum_{f\in F_{R}(n)}\frac{|f|-2}{|f|}\pi\Big\}\;\geq\;\pi.$
Then the walk $(v_{0},v_{1},\dots)$ is self-avoiding.
Proof.
Assume the walk is not self-avoiding. Let $j$ be the smallest index such that $v_{j}$ has appeared before, and let $i<j$ satisfy $v_{i}=v_{j}$. By minimality of $j$, the vertices $v_{0},\dots,v_{j-1}$ are pairwise distinct, hence
$C:=(v_{i},v_{i+1},\dots,v_{j-1},v_{j}=v_{i})$
is a simple cycle. Let $F_{C}$ be the set of faces in the bounded region enclosed by $C$.
Orient $C$ according to the traversal $v_{i}\to v_{i+1}\to\cdots\to v_{j-1}\to v_{i}$. Fix $k$ with $i<k<j$ and set $z=v_{k}$. At time $k$ the walk enters $z$ through $\langle v_{k-1},z\rangle$ and leaves through $\langle z,v_{k+1}\rangle$, which are precisely the two edges of $C$ incident to $z$. Therefore the faces of $F_{C}$ incident to $z$ are exactly the faces lying on the interior side of $C$ at $z$, hence they coincide with either $F_{L}(k)$ or $F_{R}(k)$. Using (3.4) we obtain
$\sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi\;\geq\;\pi\qquad\text{for every }z\in(C\cap V)\setminus\{v_{i}\}.$
Summing over all vertices of $C$ except $v_{i}$ yields
(3.5) $\sum_{z\in(C\cap V)\setminus\{v_{i}\}}\ \ \sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi\;\geq\;(|C|-1)\pi.$
On the other hand, by Lemma 3.2 (inequality (3.2)) applied to the cycle $C$ and the assumption $\kappa\leq 0$,
$\sum_{z\in C\cap V}\sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi\;\leq\;(|C|-2)\pi.$
Dropping the (nonnegative) contribution of the single vertex $v_{i}$ preserves the inequality, hence
$\sum_{z\in(C\cap V)\setminus\{v_{i}\}}\sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi\;\leq\;(|C|-2)\pi,$
which contradicts (3.5) since $(|C|-1)\pi>(|C|-2)\pi$. Therefore the walk must be self-avoiding. ∎
Lemma 3.5.
Assume every (finite) face has degree at least $k\geq 3$. If $\deg(v)\geq d$ and $d\cdot\frac{k-2}{k}\geq 2$, then $\kappa(v)\leq 0$. In particular, (i) if $\deg(v)\geq 6$ and $|f|\geq 3$ for all faces $f\sim v$, then $\kappa(v)\leq 0$; (ii) if $\deg(v)\geq 4$ and $|f|\geq 4$ for all faces $f\sim v$, then $\kappa(v)\leq 0$; (iii) if $\deg(v)\geq 7$ and $|f|\geq 3$ for all faces $f\sim v$, then $\kappa(v)<0$.
Lemma 3.4 has the following straightforward corollaries.
Corollary 3.6.
Let $G=(V,E)$ be a planar graph properly embedded into ${\mathbb{R}}^{2}$ such that each vertex degree is at least 6 and each face degree is at least 3. For each $v\in V$, label the incident edges of $v$ by 0, 1,…, $\mathrm{deg}(v)-1$ in counterclockwise order. Consider the following walk on $G$:
  • For any $1\leq n\leq\mathrm{length\ of\ the\ walk}-1$, if the walk visits vertex $v_{n}$ at the $n$th step, and the edge visited immediately before $v_{n}$ is the edge labelled by $a$, then the edge visited immediately after $v_{n}$ is the edge labelled by $(a+3)\mod\mathrm{deg}(v)]$.
Then the walk is self-avoiding.
Similarly, the following walk is also self-avoiding:
  • For any $1\leq n\leq\mathrm{length\ of\ the\ walk}-1$, if the walk visits vertex $v_{n}$ at the $n$th step, and the edge visited immediately before $v_{n}$ is the edge labelled by $a$, then the edge visited immediately after $v_{n}$ is the edge labelled by $(a-3)\mod\mathrm{deg}(v)]$.
Corollary 3.7.
Let $G=(V,E)$ be a planar graph, properly embedded into ${\mathbb{R}}^{2}$ such that each vertex degree is at least 4 and each face degree is at least 4. For each $v\in V$, label the incident edges of $v$ by 0, 1,…, $\mathrm{deg}(v)-1$ in counterclockwise order. Consider the following walk on $G$:
  • For any $1\leq n\leq\mathrm{length\ of\ the\ walk}-1$, if the walk visits vertex $v_{n}$ at the $n$th step, and the edge visited immediately before $v_{n}$ is the edge labelled by $a$, then the edge visited immediately after $v_{n}$ is the edge labelled by $(a+2)\mod\mathrm{deg}(v)]$.
Then the walk is self-avoiding.
Lemma 3.8.
Let $G=(V,E)$ be an infinite, connected planar graph properly embedded into ${\mathbb{R}}^{2}$ such that the minimal vertex degree is at least $7$. Then every cycle consisting of $3$ edges bounds a degree-$3$ face.
Proof.
Let $C$ be a $3$-cycle with vertices $a,b,c$. By Lemma 3.5(iii), we have $\kappa(v)\leq 0$ for all $v\in V$.
Let $F_{C}$ be the set of faces in the bounded region enclosed by $C$. Assume for contradiction that the bounded region enclosed by $C$ contains a vertex or an edge of $G$. Then at least one of $a,b,c$ is incident to at least two faces in $F_{C}$, while each of the other two vertices is incident to at least one face in $F_{C}$.
Since every finite face has degree at least $3$, we have $\frac{|f|-2}{|f|}\pi\geq\frac{\pi}{3}$. Therefore
$\sum_{z\in C\cap V}\sum_{f\in F_{C}:f\sim z}\frac{|f|-2}{|f|}\pi\;\geq\;\frac{2\pi}{3}+\frac{\pi}{3}+\frac{\pi}{3}\;=\;\frac{4\pi}{3}\;>\;\pi\;=\;(|C|-2)\pi,$
which contradicts (3.2) applied to $C$. Hence the bounded region enclosed by $C$ contains no vertices or edges of $G$, so $C$ is the boundary of a face, and that face has degree $3$. ∎
In the discussions above, we do not exclude the case that the boundary of a finite face is a closed walk visiting one vertex multiple times. The following lemma excludes the case under further assumptions on minimal vertex and face degrees.
Lemma 3.9.
Let $G=(V,E)$ be an infinite, connected planar graph properly embedded into ${\mathbb{R}}^{2}$ such that $\kappa(v)\leq 0$ for all $v\in V$. Then the boundary of every finite face is a cycle.
Proof.
Let $f$ be a finite face of $G$. Let $U_{f}$ be the unbounded component of ${\mathbb{R}}^{2}\setminus\partial f$ and let $C:=\partial U_{f}$, which is a cycle. If $\partial f\neq C$, then $|f|>|C|$.
Let $F_{C}$ be the set of faces in the bounded region enclosed by $C$. Since every edge of $C$ lies on the boundary of $f$, the face $f$ contributes at least once at each vertex of $C$ (with multiplicity in the sense of Definition 3.1), hence
$\sum_{z\in C\cap V}\sum_{g\in F_{C}:g\sim z}\frac{|g|-2}{|g|}\pi\;\geq\;|C|\cdot\frac{|f|-2}{|f|}\pi.$
Because the function $x\mapsto\frac{x-2}{x}$ is increasing for $x>0$ and $|f|>|C|$, we have
$|C|\cdot\frac{|f|-2}{|f|}\pi\;>\;|C|\cdot\frac{|C|-2}{|C|}\pi\;=\;(|C|-2)\pi,$
which contradicts (3.2) applied to $C$. Therefore $\partial f=C$, i.e. the boundary of $f$ is a cycle. ∎
Under the assumptions of Lemma 3.9, the degree $|f|$ of a finite face $f$ is equal to the number of vertices on the boundary of the face which is the same as the number of edges on the boundary of the face. We shall next construct embedded trees on graphs satisfying the assumptions of Lemma 3.9, which is crucial to the proof of Conjecture 1.1.
Lemma 3.10.
Let $G=(V,E)$ be an infinite, connected, planar graph, properly embedded into ${\mathbb{R}}^{2}$ such that the minimal vertex degree is at least 7. Then at any vertex $v\in V$ there exists a tree $T=(V_{T},E_{T})$ rooted at $v$ and embedded into $G$ such that
  • the root vertex of $T$ has degree 2; all the other vertices of $T$ have degree 3 or 4;
  • For each $\sigma\in V_{T}$, let $U_{v,\sigma}$ be all the vertices in the components of $T\setminus\{\sigma\}$ that do not contain the root $v$. Let $T_{v,\sigma}$ be the subgraph of $T$ induced by $\{\sigma\}\cup U_{v,\sigma}$. Then for each $n\geq 1$, among all the vertices of $T_{v,\sigma}$ with distance $n$ to $\sigma$, all of them have degree at least 3 and at least $\frac{1}{3}$ of them have degree 4.
  • $V_{T}\subset V$ and $E_{T}\subset{E}$.
See Figure 3.1 for an example of a tree embedded into the degree 7 triangular tilings of the hyperbolic plane (i.e., a vertex transitive graph $G$ drawn in ${\mathbb{H}}^{2}$ such that each vertex has degree 7 and each face has degree 3) satisfying the conditions of Lemma 3.10.
Figure 3.1. Tree embedding in a degree-7 triangular tiling of the hyperbolic plane: the degree-7 triangular tiling is represented by black lines, and the embedded tree is represented by red lines.
Proof.
Let $G$ be a graph satisfying condition (1) of Lemma 3.10. We will find a tree as a subgraph of $G$ recursively.
Let $v\in V$. Let $v_{0}$, $v_{1}$ be two vertices adjacent to $v$ in $G$ such that $v$, $v_{0}$, $v_{1}$ share a face. Starting from $v,v_{0}$ construct a walk
$\pi_{0}:=v,v_{0},v_{00},v_{000},\ldots,$
Starting from $v,v_{1}$ construct a walk
$\pi_{1}:=v,v_{1},v_{11},v_{111},\ldots,$
such that
  • moving along $v_{0},v,v_{1}$ in order, the face shared by $v_{0},v,v_{1}$ is on the right; and
  • moving along the walk $\pi_{0}$ starting from $v$, at each vertex $v_{0^{k}}$ ($k\geq 1$), there are exactly 3 incident faces on the right of $\pi_{0}$; and
  • moving along the walk $\pi_{1}$ starting from $v$, at each vertex $v_{1^{k}}$ ($k\geq 1$), there are exactly 3 incident faces on the left of $\pi_{1}$.
By the assumption that each vertex has degree at least 7 and Corollary 3.6, both $\pi_{0}$ and $\pi_{1}$ are infinite and self-avoiding.
Let
$\pi_{0,0}:$ $=\pi_{0}\setminus\{v\}=v_{0},v_{00},v_{000},\ldots$
$\pi_{1,1}:$ $=\pi_{1}\setminus\{v\}=v_{1},v_{11},v_{111},\ldots$
There exists $v_{01}\in V$ such that
  • $v_{01}$ is adjacent to $v_{0}$; and
  • $v_{0},v_{00},v_{01}$ share a face on the left of the walk $\pi_{0}$.
Similarly, there exist $v_{10},v_{1,\frac{1}{2}}\in V$ such that
  • both $v_{10}$ and $v_{1,\frac{1}{2}}$ are adjacent to $v_{1}$; and
  • $v_{1},v_{1,\frac{1}{2}},v_{11}$ share a face on the right of the walk $\pi_{1}$; and
  • $v_{1},v_{10},v_{1,\frac{1}{2}}$ share a face; moving along $v_{10},v_{1},v_{1,\frac{1}{2}}$ in order, the face is on the right.
Note that $v_{01}\neq v$ and $v_{10}\neq v$, $v_{1,\frac{1}{2}}\neq v$ since each vertex in $G$ has degree at least 7.
Starting from $v_{0},v_{01}$, construct a walk
$\pi_{01}:=v_{0},v_{01},v_{011},v_{0111},\ldots$
Starting from $v_{1},v_{10}$, construct a walk
$\pi_{10}:=v_{1},v_{10},v_{100},v_{1000},\ldots$
Assume that
  • moving along $v_{00},v_{0},v_{01}$ in order, the face shared by $v_{00},v_{0},v_{01}$ is on the right; and
  • moving along the walk $\pi_{01}$ starting from $v_{0}$, at each vertex $v_{01^{k}}$ ($k\geq 1$), there are exactly 3 incident faces on the left of $\pi_{01}$; and
  • moving along the walk $\pi_{10}$ starting from $v$, at each vertex $v_{10^{k}}$ ($k\geq 1$), there are exactly 3 incident faces on the right of $\pi_{10}$.
By Corollary 3.6, both walks are infinite and self-avoiding. Furthermore, let
$\tilde{\pi}_{01}:=v,\pi_{01};\qquad\tilde{\pi}_{10}:=v,\pi_{10}$
By Corollary 3.6, $\tilde{\pi}_{01}$ is self-avoiding.
We claim that $\tilde{\pi}_{10}$ is self-avoiding. Assume $\tilde{\pi}_{10}$ is not self-avoiding; we shall obtain a contradiction. Since $\pi_{10}$ is self-avoiding, if $\tilde{\pi}_{10}$ is not self-avoiding, we can find a cycle $P_{0}$ consisting of vertices of $\tilde{\pi}_{10}$ including $v$. Assume the cycle $P_{0}$ has exactly $m$ vertices denoted by $w_{1},\ldots,w_{m}$; then we must have (3.2) holds with $C$ replaced by $P_{0}$, $n$ replaced by $m$ and $V_{i}$ replaced by $w_{i}$.
Under the assumption that each vertex has degree at least 7 and each face has degree at least 3, $\kappa(z)<0$; and
(3.6) $\frac{|f|-2}{|f|}\geq\frac{1}{3}$
Note that each vertex along $P_{0}$ except $v$ and $v_{1}$ is incident to at least 3 faces in the bounded region $R_{P_{0}}$ enclosed $P_{0}$; $v$ is incident to at least 1 face in $R_{P_{0}}$ and $v_{1}$ is incident to at least 2 faces in $R_{P_{0}}$. Then we have
$\sum_{z\in V\cap P_{0}}\sum_{f\in F_{P_{0}}:f\sim z}\frac{|f|-2}{|f|}\pi\geq(m-2)\pi+\sum_{f\in F_{P_{0}}:f\sim v,\mathrm{or}\ f\sim v_{1}}\frac{|f|-2}{|f|}\pi\geq(m-2)\pi+\pi>(m-2)\pi,$
which contradicts (3.2), and therefore $\tilde{\pi}_{10}$ is self-avoiding.
We claim that $\pi_{01}$ and $\pi_{10}$ never intersect each other. Otherwise let $w\in V\cap\pi_{01}\cap\pi_{10}$ be an intersection vertex of $\pi_{01}$ and $\pi_{10}$, such that the portion of $\pi_{01}$ between $v_{0}$ and $w$, the portion of $\pi_{10}$ between $v_{1}$ and $w$ and the edges $v_{0}v$, $v_{1}v$ form a cycle $P$ in the plane. Assume the cycle $P$ has exactly $n$ vertices; then we must have (3.2) holds. Under the assumption that each vertex has degree at least 7 and each face has degree at least 3, $\kappa(z)<0$; we have (3.6).
The following cases might occur
  1. (a)
    $w\neq v_{0}$ and $w\neq v_{1}$. Under the assumption that $v_{0}$ and $v_{1}$ have degrees at least 7, $v_{0}$ is incident to at least 3 faces in $F_{P}$ and $v_{1}$ is incident to at least 2 faces in $F_{P}$; where $F_{P}$ is the set of faces in the bounded region enclosed by $P$.
    Note that each vertex along $P$ except $v$, $w$ and $v_{1}$ are incident to at least 3 faces in $R_{P}$; $v$ and $w$ are incident to at least 1 face in $R_{P}$ and $v_{1}$ is incident to at least 2 faces in $R_{P}$. Then we have
    (3.7) $\sum_{z\in V\cap P}\sum_{f\in F_{P}:f\sim z}\frac{|f|-2}{|f|}\pi\geq(n-3)\pi+\sum_{f\in F_{P}:f\sim v,\mathrm{or}\ f\sim w,\ \mathrm{or}\ f\sim v_{1}}\frac{|f|-2}{|f|}\pi\geq(n-3)\pi+\frac{4\pi}{3}>(n-2)\pi.$
  2. (b)
    $w=v_{0}$. (3.7) still holds.
  3. (c)
    $w=v_{1}$. Each vertex along $P$ except $v$, $w$ are incident to at least 3 faces in $R_{P}$; $v$ and $w$ are incident to at least 1 face in $F_{P}$ and $v_{1}$.Then we have
    $\sum_{z\in V\cap P}\sum_{f\in F_{P}:f\sim z}\frac{|f|-2}{|f|}\pi\geq(n-2)\pi+\sum_{f\in F_{P}:f\sim v,\mathrm{or}\ f\sim w,\ }\frac{|f|-2}{|f|}\pi\geq(n-2)\pi+\frac{2\pi}{3}>(n-2)\pi.$
Hence (3.1) never holds, and therefore $\pi_{01}$ and $\pi_{10}$ are disjoint.
We repeat the same construction with $(v_{0},v,v_{1})$ replaced by $(v_{00},v_{0},v_{01})$.
Starting from $v_{1},v_{1,\frac{1}{2}}$, we construct two walks
$\pi_{1,\frac{1}{2},0}:=v_{1},v_{1,\frac{1}{2}},v_{1,\frac{1}{2},0},v_{1,\frac{1}{2},0,0},\ldots$
$\pi_{1,\frac{1}{2},1}:=v_{1},v_{1,\frac{1}{2}},v_{1,\frac{1}{2},1},v_{1,\frac{1}{2},1,1},\ldots$
such that
  • moving along the walk $\pi_{1,\frac{1}{2},0}$ staring from $v_{1}$, at each vertex $\pi_{1,\frac{1}{2},0^{k}}$ ($k\geq 0$), there are exactly 3 incident faces on the right.
  • moving along the walk $\pi_{1,\frac{1}{2},1}$ staring from $v_{1}$, at each vertex $\pi_{1,\frac{1}{2},1^{k}}$ ($k\geq 0$), there are exactly 3 incident faces on the left.
Let
$\tilde{\pi}_{1,\frac{1}{2},1}:=\{v\}\cup\pi_{1,\frac{1}{2},1};\qquad\tilde{\pi}_{1,\frac{1}{2},0}:=\{v\}\cup\pi_{1,\frac{1}{2},0}$
By Corollary 3.6, both $\tilde{\pi}_{1,\frac{1}{2},1}$ and $\tilde{\pi}_{1,\frac{1}{2},0}$ are infinite and self-avoiding. Moreover, using Lemma 3.4, one can prove that
  1. (A)
    The intersection of any two paths in $\pi_{1},\pi_{10},\pi_{1,\frac{1}{2},1},\pi_{1,\frac{1}{2},0}$ is $\{v_{1}\}$.
  2. (B)
    $\left[\pi_{1,\frac{1}{2},0}\cup\pi_{1,\frac{1}{2},1}\right]\cap\pi_{0}=\emptyset$ and $\left[\pi_{1,\frac{1}{2},0}\cup\pi_{1,\frac{1}{2},1}\right]\cap\pi_{01}=\emptyset$
Let $v$ be the level-0 vertex, $v_{0},v_{1}$ be level-1 vertices, and $v_{00},v_{01},v_{10},v_{1,\frac{1}{2}},v_{11}$ be the level-2 vertices. In general For $k\geq 2$, define the set $S_{k}$ of level-$k$ vertices as follows
(3.8) $S_{k}:=\left\{v_{b}:b=(b_{1},\ldots,b_{k})\in\left\{0,\frac{1}{2},1\right\}^{k};\mathrm{if}\ b_{j}=\frac{1}{2},\ \mathrm{then}\ j\geq 2,\ \mathrm{and}\ b_{j-1}=1.\right\}.$
Assume we defined all the level-$k$ vertices. For each $v_{b}\in S_{k}$, the following cases might occur
  • $b_{k}=0$: in this case we define 2 paths $\pi_{b,0}$, $\pi_{b,1}$ as defining $\pi_{00}$ and $\pi_{01}$ with $(v,v_{0})$ replaced by $(v_{b_{1},\ldots,b_{k-1}},v_{b})$.
  • $b_{k}=1$: in this case we define 3 paths $\pi_{b,0}$, $\pi_{b,\frac{1}{2}}$ $\pi_{b,1}$ as defining $\pi_{10}$ $\pi_{1,\frac{1}{2}}$ and $\pi_{11}$ with $(v,v_{1})$ replaced by $(v_{b_{1},\ldots,b_{k-1}},v_{b})$.
  • $b_{k}=\frac{1}{2}$: in this case we define 2 paths $\pi_{b,0}$, $\pi_{b,1}$ as defining $\pi_{00}$ and $\pi_{01}$ with $(v,v_{0})$ replaced by $(v_{b_{1},\ldots,b_{k-1}},v_{b})$.
This way we find a tree $T$ whose vertex set consists of $\{v,v_{0},v_{1}\}\cup_{k\geq 2}S_{k}$ and edge set consists of all the edges along a path $\pi_{b}$ such that for some $k\geq 1$ $b=(b_{1},\ldots,b_{k})\in\left\{0,\frac{1}{2},1\right\}^{k};\mathrm{if}\ b_{j}=\frac{1}{2},\ \mathrm{then}\ j\geq 2,\ \mathrm{and}\ b_{j-1}=1$ as a subgraph of $G$. Then Part (1) of Lemma 3.10 follows. ∎
Proposition 3.11.
Let $G=(V,E)$ be an infinite, connected planar graph properly embedded in $\mathbb{R}^{2}$ with minimal vertex degree at least 7. Let $T\subseteq G$ be the embedded rooted tree given by Lemma 3.10. Set $p_{T}:=p_{c}^{\mathrm{site}}(T)$. Then for every
$p\in(p_{T},\ 1-p_{T}),$
${\mathbb{P}}_{p}$-a.s. there exist infinitely many infinite $1$-clusters and infinitely many infinite 0-clusters in $G$.
Proof.
Step 1: $p_{T}<\tfrac{1}{2}$. By [21, Thm. 6.2] (see also [22, Thm. 5.15]), for any infinite locally finite tree,
$p_{c}^{\mathrm{site}}(T)=\frac{1}{\mathrm{br}(T)},$
where $\mathrm{br}(T)$ denotes the branching number. Lemma 3.10(2) implies a uniform forward expansion: if $M_{n}^{\sigma}$ is the number of descendants of $\sigma$ at distance $n$ in the forward subtree $T_{v,\sigma}$, then
(3.9) $M_{n}^{\sigma}\ \geq\ 2\Bigl(\frac{7}{3}\Bigr)^{n-1}\qquad(n\geq 1),$
and consequently $\mathrm{br}(T)\geq\frac{7}{3}$. Hence
$p_{T}=\frac{1}{\mathrm{br}(T)}\leq\frac{3}{7}<\frac{1}{2}.$
Fix now $p\in(p_{T},1-p_{T})$ and set q:=1pq:=1-p. Then $p>p_{T}$ and $q>p_{T}$.
Step 2: A local separation event. Fix $p\in\bigl(p_{c}^{\mathrm{site}}(T),\,1-p_{c}^{\mathrm{site}}(T)\bigr)$ and set q:=1pq:=1-p. Recall that each vertex in $T$ can be represented by $v_{r}$ with $r\in\{0,\frac{1}{2},1\}^{<\infty}$. For each $r$, let $T(r)\subset T$ denote the forward subtree of $T$ rooted at $v_{r}$.
Define the event $A_{r}$ that
$\eta(v_{r})=\eta(v_{r0})=\eta(v_{r1})=\eta(v_{r00})=\eta(v_{r10})=0,$
and
$v_{r00}\stackrel{{\scriptstyle 0}}{{\longleftrightarrow}}\infty\ \text{in}\ T(r00),\qquad v_{r10}\stackrel{{\scriptstyle 0}}{{\longleftrightarrow}}\infty\ \text{in}\ T(r10),$
and
$v_{r01}\stackrel{{\scriptstyle 1}}{{\longleftrightarrow}}\infty\ \text{in}\ T(r01),\qquad v_{r11}\stackrel{{\scriptstyle 1}}{{\longleftrightarrow}}\infty\ \text{in}\ T(r11).$
Since $p>p_{c}^{\mathrm{site}}(T)$ and $q>p_{c}^{\mathrm{site}}(T)$, each of the four percolation events has positive probability in the corresponding forward subtree, and these subtrees are pairwise disjoint. Therefore
${\mathbb{P}}_{p}(A_{r})\geq a(p)>0\qquad\text{for all }r\in\{0,1\}^{<\infty}.$
On $A_{r}$, the closed path
$v_{r00}-v_{r0}-v_{r}-v_{r1}-v_{r10}$
together with the two infinite closed rays inside $T(r00)$ and $T(r10)$ yields an infinite closed cluster $\Xi_{r}$ in $G$ whose embedding contains a doubly-infinite simple curve. By the left/right embedded-tree construction (Lemma 3.10 and the planar disjointness properties used there), the two forward subtrees $T(r01)$ and $T(r11)$ lie in different unbounded components of ${\mathbb{R}}^{2}\setminus\Xi_{r}$. Consequently, any $1$-path in $G$ connecting the $1$-clusters intersecting $T(r01)$ and $T(r11)$ would have to cross $\Xi_{r}$, which is impossible. Hence, on $A_{r}$ there are at least two distinct infinite $1$-clusters in $G$. Moreover, $\Xi_{r}$ is an infinite 0-cluster in $G$.
Step 3: Borel–Cantelli. Choose an infinite sequence of pairwise incomparable words $(r_{m})_{m\geq 1}$ (e.g. $r_{m}:=0\left(1,\frac{1}{2}\right)^{m}$). Then the vertex sets involved in the events $A_{r_{m}}$ are disjoint, so $(A_{r_{m}})_{m\geq 1}$ are independent and
$\sum_{m=1}^{\infty}{\mathbb{P}}_{p}(A_{r_{m}})\geq\sum_{m=1}^{\infty}a(p)=\infty.$
By the second Borel–Cantelli lemma, $A_{r_{m}}$ occurs for infinitely many $m$ almost surely. Each occurrence produces (inside the disjoint region below $v_{r_{m}}$) at least one more infinite $1$-clusters ${\mathbb{P}}_{p}$-a.s. $G$ has infinitely many infinite $1$-clusters, and infinitely many infinite 0-clusters by symmetry. ∎
Figure 3.2. Infinite open clusters in the tree rooted at $v_{r01}$ is separated from the infinite open clusters in the tree rooted at $v_{r11}$ by the infinite closed cluster occupying $v_{r}$, $v_{r0},v_{r1}$, $v_{r00},v_{r10}$
Remark 3.12.
If we replace the assumption ‘the minimal vertex degree is at least 7” in Lemmas 3.10 and 3.11 by the minimal vertex degree is at least 5; and the minimal face degree is at least 4”, the same conclusion can be proved similarly.
4. Characterization of Critical Percolation Probability
In a powerful refinement of a classical argument of Hammersley [15], Duminil-Copin and Tassion [9] showed, for transitive $G$, that the critical value $p_{c}(G)$ may be characterized in terms of the mean number of points on the surface of a box that are connected to its root. This work is extended in this section to general locally finite graphs without the transitive or quasi-transitive assumptions. The proof utilizes the technique of differential inequalities ([1]). The techniques developed in this section can also be applied to [19] to establish a vertex-cut characterization of $p_{c}^{site}$ due to Kahn ([17]), and to disprove an edge-cut characterization of $p_{c}^{site}$ proposed by Lyons and Peres ([22]). Both vertex-cut and edge-cut characterizations for $p_{c}^{bond}$ and for $p_{c}^{site}$ on bounded-degree graphs were previously proved in [28]. The main advantage of the techniques presented here is that they remove the need for the bounded-degree assumption.
Let $G=(V,E)$ be a graph. For each $p\in(0,1)$, let $\mathbb{P}_{p}$ be the probability measure of the i.i.d. Bernoulli($p$) site percolation on $G$. For each $S\subset V$, let $S^{\circ}$ consist of all the interior vertices of $S$, i.e., vertices all of whose neighbors are in $S$ as well. For each $S\subseteq V$, $v\in S$, define
[…]
where
[…]
  • $\partial_{V}y$ consists of all the vertices adjacent to $y$.
The following technical lemmas (Lemmas 4.1 and 4.2) were proved in [19], based on an adaptation of arguments in [9].
Lemma 4.1.
Let $G=(V,E)$ be an infinite, connected, locally finite graph. The critical site percolation probability on $G$ is given by
$\tilde{p}_{c}=\sup\{p\geq 0:\exists\epsilon_{0}>0,\mathrm{s.t.}\forall v\in V,\exists S_{v}\subseteq V\ \mathrm{satisfying}\ |S_{v}|<\infty\ \mathrm{and}\ v\in S_{v}^{\circ},\varphi_{p}^{v}(S_{v})\leq 1-\epsilon_{0}\}$
Moreover,
  1. (1)
    If $p>\tilde{p}_{c}$, a.s. there exists an infinite 1-cluster; moreover, for any $\epsilon>0$ there exists a vertex $w$, such that
    (4.1) $\forall S_{w}\subseteq V\ \text{finite with }w\in S_{w}^{\circ},\quad\varphi_{q}^{w}(S_{w})>1-\epsilon_{1}\qquad\forall q\geq p_{1}.;\ \forall q\geq p_{1}$
    where $p_{1},\epsilon_{1}$ are such that
    (4.2) $p_{1}\in(\tilde{p}_{c},p);\ \epsilon_{1}\in(0,\epsilon);\ \left(\frac{1-p}{1-p_{1}}\right)^{1-\epsilon_{1}}<\left(\frac{1-p}{1-\tilde{p}_{c}}\right)^{1-\epsilon}.$
    Any vertex $w$ satisfying (4.1) also satisfies
    (4.3) $\mathbb{P}_{p}(w\leftrightarrow\infty)\geq 1-\left(\frac{1-p}{1-\tilde{p}_{c}}\right)^{1-\epsilon}$
  2. (2)
    If $p<\tilde{p}_{c}$, then for any vertex $v\in V$
    (4.4) $\mathbb{P}_{p}(v\leftrightarrow\infty)=0.$
In particular, (1) and (2) implies that $p_{c}^{site}(G)=\tilde{p}_{c}$
Lemma 4.2.
Let $G=(V,E)$ be an infinite, connected, locally finite graph. Let $p>0$, $u\in S\subset A$ and $B\cap S=\emptyset$. Then
  • If $u\in S^{\circ}$
[…]
  • If $u\in S\setminus S^{\circ}$,
[…]
Lemma 4.3.
Let $G=(V,E)$ be an infinite, connected, locally finite graph. For each $p>p_{c}^{site}(G)$ and $\epsilon>0$, there exist infinitely many vertices in $V$ satisfying (4.3).
Proof.
Fix $p>p_{c}^{site}(G)$ and $\epsilon>0$. Choose $p_{1}\in(p_{c}^{site}(G),p)$ and $\epsilon_{1}\in(0,\epsilon)$ as in Lemma 4.1(1), so that (4.2) holds. Set $\delta:=\epsilon_{1}/2$. Define
(4.5) $V_{p_{1},\delta}:=\Bigl\{v\in V:\ \forall\,S\subset V\ \text{finite with }v\in S^{\circ},\ \varphi_{p_{1}}^{v}(S)\geq 1-\delta\Bigr\}.$
Step 1: $|V_{p_{1},\delta}|=\infty$. Recall $V_{p_{1},\delta}$ defined in (4.5). Assume for contradiction that $|V_{p_{1},\delta}|<\infty$. Then there exist $v_{0}\in V$ and $N\in\mathbb{N}$ such that $V_{p_{1},\delta}\subseteq B(v_{0},N)$. Let
[…]
For each $x\in W$, since $x\notin V_{p_{1},\delta}$, there exists a finite set $S_{x}\subseteq W$ with $x\in S_{x}^{\circ}$ and $\varphi_{p_{1}}^{x}(S_{x})\leq 1-\delta$. Applying Lemma 4.2 with $A=W$, $B=\infty$ and $S=S_{x}$, we obtain
[…]
$\leq\varphi_{p_{1}}^{x}(S_{x})\,M\leq(1-\delta)M.$
[…]
Step 2: infinitely many vertices satisfy (4.3). Pick infinitely many vertices $w\in V_{p_{1},\delta}$. Then for every such $w$, every finite $S_{w}\subset V$ with $w\in S_{w}^{\circ}$, and every $q\geq p_{1}$, monotonicity in $p$ gives
$\varphi_{q}^{w}(S_{w})\ \geq\ \varphi_{p_{1}}^{w}(S_{w})\ \geq\ 1-\delta\ >\ 1-\epsilon_{1},$
so $w$ satisfies (4.1). By Lemma 4.1(1), any vertex satisfying (4.1) also satisfies (4.3) at parameter $p$. Since there are infinitely many such $w$, the lemma follows. ∎
Proof of Proposition 1.3. For each $v\in V$, let $C(v)$ be the 1-cluster including $v$. If $v$ is closed, then $C(v)=\emptyset$. We have
$\mathbb{P}_{p}(u\leftrightarrow v)\geq\mathbb{P}_{p}(u\leftrightarrow\infty,v\leftrightarrow\infty,\mathcal{A}_{1})=\mathbb{P}_{p}(v\leftrightarrow\infty,\mathcal{A}_{1})-\mathbb{P}_{p}(v\leftrightarrow\infty,\mathcal{A}_{1},u\nleftrightarrow\infty)$
Note that
$\mathbb{P}_{p}(v\leftrightarrow\infty,\mathcal{A}_{1},u\nleftrightarrow\infty)$ $=\sum_{S:[u\in S,|S|<\infty]\ \mathrm{or}\ S=\emptyset}\mathbb{P}_{p}(v\leftrightarrow\infty,\mathcal{A}_{1}|C(u)=S)\mathbb{P}_{p}(C(u)=S)$
$\leq\mathbb{P}_{p}(v\leftrightarrow\infty,\mathcal{A}_{1})\sum_{S:[u\in S,|S|<\infty]\ \mathrm{or}\ S=\emptyset}\mathbb{P}_{p}(C(u)=S)$
$=\mathbb{P}_{p}(v\leftrightarrow\infty,\mathcal{A}_{1})\mathbb{P}_{p}(u\nleftrightarrow\infty)$
Hence we have
$\mathbb{P}_{p}(u\leftrightarrow\infty,v\leftrightarrow\infty,\mathcal{A}_{1})\geq\mathbb{P}_{p}(v\leftrightarrow\infty,\mathcal{A}_{1})\mathbb{P}_{p}(u\leftrightarrow\infty)$
Similarly we have
$\mathbb{P}_{p}(v\leftrightarrow\infty,\mathcal{A}_{1})\geq\mathbb{P}_{p}(v\leftrightarrow\infty)\mathbb{P}_{p}(\mathcal{A}_{1})$
[…]
Proposition 1.3 may also be proved using the van den Berg–Kesten–Reiner (BKR) inequality (see [29, 26]; see also [6, 25]).
Proof of Corollary 1.4. Let $\mathcal{A}_{f}$ be the event that the number of infinite 1-clusters is finite and nonzero. Let $p>p_{c}^{site}(G)$. It suffices to show that if $\mathbb{P}_{p}(\mathcal{A}_{f})>0$, then $p\geq p_{conn}$.
If $\mathbb{P}_{p}(\mathcal{A}_{f})>$, then $\mathbb{P}_{p}(\mathcal{A}_{1})>0$. Let $u,v\in V_{p,\epsilon}$ for some $\epsilon>0$. We have
Given $u,v\in V_{p,\epsilon}$, by Lemma 1.3 we obtain
$\mathbb{P}_{p}(u\leftrightarrow v)\geq\left[1-\left(\frac{1-p}{1-p_{c}}\right)^{1-\epsilon}\right]^{2}\mathbb{P}_{p}(\mathcal{A}_{1})>0.$
[…]
5. Embedded Forests
In this section we prove exponential decay of point-to-point connection probabilities in the matching graph $G_{*}$ (and hence also in $G$) whenever $p<1-p_{c}^{\mathrm{site}}(T)$. The key is to construct, along a fixed geodesic $l_{uw}$ between two vertices $u$ and $w$, an embedded forest whose components separate $u$ from $w$ more and more often as $d(u,w)\to\infty$. This construction uses the planarity of $G$ and the assumption that the minimum vertex degree is at least $7$.
Definition 5.1 (Chandeliers and anti-chandeliers).
Let $G=(V,E)$ be an infinite, connected graph properly embedded in $\mathbb{R}^{2}$, with minimum vertex degree at least $7$. Fix $v\in V$. A chandelier $R_{v}$ rooted at $v$ is a tree embedded in $G$ together with distinguished vertices $v_{1},v_{2}\in V(R_{v})$ such that:
  1. (C1)
    $v_{1}$ is the unique neighbor of $v$ in $R_{v}$, and $v_{2}$ is the unique neighbor of $v_{1}$ in $R_{v}$ distinct from $v$ (so the path from $v$ to $v_{2}$ in $R_{v}$ is $v,v_{1},v_{2}$).
  2. (C2)
    Traversing the path $v,v_{1},v_{2}$ in this order, the vertex $v_{1}$ has exactly three incident faces on the left.
  3. (C3)
    The subtree of $R_{v}$ rooted at $v_{2}$, denoted by $T_{v_{2}}$, is isomorphic to the tree constructed in Lemma 3.10(1).
  4. (C4)
    Let $L_{1}$ and $L_{2}$ be the two boundary rays of $T_{v_{2}}$ with $L_{1}\cap L_{2}=\{v_{2}\}$, and let $Q\subset\mathbb{R}^{2}$ be the (closed) region bounded by $L_{1}\cup L_{2}$ that contains the embedding of $T_{v_{2}}$. Then:
    • traversing $v_{1},v_{2},L_{1}$ in this order, at $v_{2}$ there are at least $3$ incident faces on the left outside $Q$, and exactly $1$ incident face on the right inside $Q$;
    • traversing $v_{1},v_{2},L_{2}$ in this order, at $v_{2}$ there are at least $3$ incident faces on the right outside $Q$, and exactly $1$ incident face on the left inside $Q$.
An anti-chandelier is defined in the same way, except that in (C2) and (C4) we interchange “left” and “right”.
If $R_{v}$ is a chandelier, we call
$\langle v,v_{1}\rangle\cup\langle v_{1},v_{2}\rangle\cup L_{1}\quad\text{(resp.\ }\langle v,v_{1}\rangle\cup\langle v_{1},v_{2}\rangle\cup L_{2}\text{)}$
the left boundary (resp. right boundary) of $R_{v}$. If $R_{v}$ is an anti-chandelier, the roles of $L_{1}$ and $L_{2}$ are swapped, so the same sets are called the right and left boundaries, respectively.
See Figure 5.1 for chandeliers and anti-chandeliers.
Figure 5.1. In the left graph, the black lines represent a chandelier, the blue lines represent incident edges on the left of a chandelier. In the right graph, the black lines represent an anti-chandelier, and blue lines represent incident edges on the right of an anti-chandelier.
For each $v\in V$, define
[…]
Here $|f|$ denotes the degree of the face $f$, i.e., the number of edges along its boundary. If all faces incident to $v$ are infinite, we set $DF_{v}=$-$\infty$.
Let $u,w\in V$ be distinct vertices, and fix a geodesic (shortest path) in $G$ from $u$ to $w$,
$l_{uw}=(z_{0},z_{1},\ldots,z_{n}),\qquad z_{0}=u,\ z_{n}=w,$
where $z_{i-1}$ and $z_{i}$ are adjacent for all $1\leq i\leq n$. For $x=z_{j}\in l_{uw}$ we call $j$ the $z$-index of $x$. We orient $l_{uw}$ from $u$ to $w$ (increasing $z$-index). All references to “left” and “right” are with respect to this orientation.
We now construct, along $l_{uw}$, a sequence of chandeliers (or empty sets)
$R_{1},R_{2},\ldots$
embedded to the left of $l_{uw}$, and a sequence of anti-chandeliers (or empty sets)
$U_{1},U_{2},\ldots$
embedded to the right of $l_{uw}$.
Left side. Let $\zeta_{1}$ be the first vertex on $l_{uw}$ (excluding $z_{0}$ and the last three vertices $z_{n-2},z_{n-1},z_{n}$) such that either
  1. (L1)
    $\zeta$ is incident to at least two faces on the left of $l_{uw}$ (with respect to the orientation $u\to w$), or
  2. (L2)
    $\zeta$ is incident to exactly one infinite face on the left of $l_{uw}$ (with respect to the orientation $u\to w$).
If (L2) holds, we set $R_{1}=\emptyset$. If (L1) holds, let $b_{1}$ be the predecessor of $\zeta_{1}$ on $l_{uw}$ and choose the incident edge $e_{1}=\langle\zeta_{1},a_{1}\rangle$ on the left of $l_{uw}$ with $a_{1}\notin l_{uw}$ such that there are no other edges of $G$ in the angle (on the left side) between $e_{1}$ and $\langle b_{1},\zeta_{1}\rangle$. Define $R_{1}$ to be the chandelier at $\zeta_{1}$ starting from $e_{1}$.
Inductively, suppose $\zeta_{s}$ and $R_{s}$ have been defined. Let $\zeta_{s+1}$ be the next vertex after $\zeta_{s}$ along $l_{uw}$ satisfying (L1) or (L2) with minimal $z$-index. If $\zeta_{s+1}$ satisfies (L2), set $R_{s+1}=\emptyset$. If $\zeta_{s+1}$ satisfies (L1), define $b_{s+1}$ and pick $e_{s+1}=\langle\zeta_{s+1},a_{s+1}\rangle$ analogously. If $R_{s}\neq\emptyset$ and $a_{s+1}=a_{s}$, set $R_{s+1}=R_{s}$; otherwise let $R_{s+1}$ be the chandelier at $\zeta_{s+1}$ starting from $e_{s+1}$.
We stop once we have passed $z_{n-1}$. Since $l_{uw}$ is finite, the construction terminates after finitely many steps.
Lemma 5.2.
If $R_{j}\neq\emptyset$ and $R_{j+3}\neq\emptyset$, then $a_{j}\neq a_{j+3}$.
Proof.
From the construction of $l_{uw}$ we see that $l_{\zeta_{j}\zeta_{j+3}}$ is the shortest path in $G$ joining $\zeta_{j}$ and $\zeta_{j+3}$ which has length at least 3. If $a_{j}=a_{j+3}$, then $\zeta_{j},a_{j}(=a_{j+3}),\zeta_{j+3}$ form a path of length 2 joining $\zeta_{j}$ and $\zeta_{j+3}$. The contradiction implies the lemma. ∎
Right side. Define the sequence $(\xi_{i},U_{i})_{i\geq 1}$ on the right of $l_{uw}$ mutatis mutandis from the above left-side construction, with the following replacements:
$\text{left}\leftrightarrow\text{right},\qquad(\zeta_{i},b_{i},e_{i},a_{i},R_{i})\leftrightarrow(\xi_{i},c_{i},f_{i},d_{i},U_{i}),\qquad\text{chandelier}\leftrightarrow\text{anti-chandelier}.$
In particular, in the inductive step the identification case is interpreted as $U_{s+1}=U_{s}$ only if $U_{s}\neq\emptyset$ and $d_{s+1}=d_{s}$; otherwise $U_{s+1}$ is defined as a new anti-chandelier (or $\emptyset$ if the infinite-face case occurs).
Lemma 5.3.
If $U_{j}\neq\emptyset$ and $U_{j+3}\neq\emptyset$, then $d_{j}\neq d_{j+3}$.
Let
$\mathcal{R}_{uw}$ $=\{R_{1},R_{2},\ldots\}$
$\mathcal{U}_{uw}$ $=\{U_{1},U_{2},\ldots\}$
More precisely, $\mathcal{R}_{uw}$ is the set of all the chandeliers and emptysets constructed on the left of $l_{uw}$ as above; and $\mathcal{U}_{uw}$ is the set of all the anti-chandeliers and emptysets constructed on the right of $l_{uw}$ as above.
Figure 5.2. Embedded forest
Fix a shortest path $l_{uw}=(z_{0}=u,z_{1},\dots,z_{n}=w)$, oriented from $u$ to $w$. For $x,y\in l_{uw}$, write $l_{uw}[x,y]$ for the (unique) subpath of $l_{uw}$ joining $x$ and $y$ (including endpoints). For a rooted chandelier/anti-chandelier $H$, write $\mathrm{rt}(H)$ for its root.
Setup 5.4.
Let $G,u,w,l_{uw},\mathcal{R}_{uw},\mathcal{U}_{uw}$ be as above, and let $\{R_{m}\}_{m\geq 1}$ (resp. $\{U_{m}\}_{m\geq 1}$) be the left (resp. right) sequence constructed along $l_{uw}$.
(Indexing of distinct nonempty objects). Let $i_{1}<i_{2}<\cdots<i_{k}$ be the indices of the first occurrences of the distinct nonempty chandeliers in $\{R_{m}\}$, i.e.
  1. (i)
    $R_{i_{s}}\neq\emptyset$ for all $s$ and $R_{i_{s}}\neq R_{i_{s^{\prime}}}$ for $s\neq s^{\prime}$;
  2. (ii)
    for every $m$ with $R_{m}\neq\emptyset$, there exists a unique $s$ such that $R_{m}=R_{i_{s}}$, and $i_{s}=\min\{m^{\prime}:\,R_{m^{\prime}}=R_{i_{s}}\}$.
Define $j_{1}<j_{2}<\cdots<j_{r}$ analogously for the distinct nonempty anti-chandeliers in $\{U_{m}\}$.
(Boundary rays and first intersection). For each $s$ (resp. $t$), let $\partial_{L}R_{i_{s}},\partial_{R}R_{i_{s}}$ (resp. $\partial_{L}U_{j_{t}},\partial_{R}U_{j_{t}}$) be the left/right boundary ray, oriented away from its root $\zeta_{i_{s}}$ (resp. $\xi_{j_{t}}$). For an oriented ray $\gamma$ and a set $A$ with $\gamma\cap A\neq\emptyset$, let $\mathrm{first}(\gamma,A)$ denote the first intersection vertex of $\gamma$ with $A$ encountered when traversing $\gamma$ from its root.
Remark 5.5.
Whenever we refer to “an edge incident to $x$ on the left of a ray”, we implicitly assume $x$ is not the root of that ray.
Lemma 5.6.
Assume Setup 5.4.
  1. (1)
    (Left-side chandeliers) For the sequence $R_{i_{1}},\dots,R_{i_{k}}$:
    1. (a)
      If $2\leq s\leq k$, then the edge immediately preceding $\mathrm{first}(\partial_{L}R_{i_{s}},l_{uw}[\zeta_{i_{s-1}},\zeta_{i_{s}}))$ along $\partial_{L}R_{i_{s}}$ cannot be on the left of $l_{uw}[\zeta_{i_{s-1}},\zeta_{i_{s}})$.
    2. (b)
      If $1\leq s\leq k-1$, then the edge immediately preceding $\mathrm{first}(\partial_{R}R_{i_{s}},l_{uw}(\zeta_{i_{s}},\zeta_{i_{s+1}}])$ along $\partial_{R}R_{i_{s}}$ cannot be on the left of $l_{uw}(\zeta_{i_{s}},\zeta_{i_{s+1}}]$.
  2. (2)
    (Right-side anti-chandeliers) For the sequence $U_{j_{1}},\dots,U_{j_{r}}$:
    1. (a)
      If $2\leq t\leq r$, then the edge immediately preceding $\mathrm{first}(\partial_{R}U_{j_{t}},l_{uw}[\xi_{j_{t-1}},\xi_{j_{t}}))$ along $\partial_{R}U_{j_{t}}$ cannot be on the right of $l_{uw}[\xi_{j_{t-1}},\xi_{j_{t}})$.
    2. (b)
      If $1\leq t\leq r-1$, then the edge immediately preceding $\mathrm{first}(\partial_{L}U_{j_{t}},l_{uw}(\xi_{j_{t}},\xi_{j_{t+1}}])$ along $\partial_{L}U_{j_{t}}$ cannot be on the right of $l_{uw}(\xi_{j_{t}},\xi_{j_{t+1}}]$.
Proof.
We prove (1a) and (1b); all other parts follow from the same argument after interchanging left and right (and replacing $R$ by $U$ when needed).
Proof of (1a). Let
$a:=\mathrm{first}(\partial_{L}R_{i_{s}},l_{uw}[\zeta_{i_{s-1}},\zeta_{i_{s}}).$
Assume for contradiction that the edge along $\partial_{L}R_{i_{s}}$ immediately preceding $a$ is on the left of $l_{uw}[\zeta_{i_{s-1}},\zeta_{i_{s}})$.
Set
$P:=l_{uw}[a,\zeta_{i_{s}}],$
so $P$ is a geodesic segment of $l_{uw}$.
Step 1: build a simple cycle. Let $\Gamma$ be the subpath of the boundary ray $\partial_{L}R_{i_{s}}$ joining $\zeta_{i_{s}}$ to $a$ (traversed from $\zeta_{i_{s}}$ towards $a$). By the choice of $a$ and the fact that boundary rays are self-avoiding, the union
$C:=P\cup\Gamma$
is a simple cycle. Let $F_{C}$ be the set of faces in the bounded component of $\mathbb{R}^{2}\setminus C$.
Step 2: a unique interior face along $P$. By the construction of the points $\zeta_{i_{s}}$ on $l_{uw}$, every internal vertex of $P$ (i.e. of $P\setminus\{\zeta_{i_{s}},a\}$) is incident to exactly one face on the left of $l_{uw}$. Since the interior of $C$ lies on the left side of $P$, it follows that the interior of $C$ is incident to every edge of $P$ through a single face, denoted $f_{0}\in F_{C}$, and $\partial f_{0}$ contains $P$ as a subwalk.
[…]
$|f_{0}|\geq 2m,$
and since every finite face has degree at least $3$, we may record
$|f_{0}|\geq\max\{2m,3\}.$
Step 3: Gauss–Bonnet contradiction. Apply Lemma 3.2 (3.2) to the simple cycle $C$:
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}\frac{|f|-2}{|f|}\pi\ \leq\ (|C|-2)\pi.$
Let $w(f):=\frac{|f|-2}{|f|}\pi$.
(i) Contribution from $P$. The face $f_{0}$ is incident to every vertex of $P$, hence for each $v\in P\cap V$,
$\sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ w(f_{0}).$
Therefore the total contribution from vertices of $P$ is at least $(m+1)w(f_{0})$.
(ii) Contribution from $\Gamma\setminus\{\zeta_{i_{s}},a\}$. Each internal vertex of $\Gamma$ lies on a boundary ray of a chandelier, so by the defining turning rule of chandeliers the interior of $C$ occupies at least three face-sectors around such a vertex. All those faces belong to $F_{C}$ and each has degree at least $3$, hence each such vertex contributes at least $3\cdot(\pi/3)=\pi$. There are $|\Gamma|-1$ internal vertices on $\Gamma$, so their total contribution is at least $(|\Gamma|-1)\pi$.
Combining (i) and (ii) yields
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ (|\Gamma|-1)\pi+(m+1)w(f_{0}).$
Since $|C|=m+|\Gamma|$, it remains to compare the right-hand side with $(|C|-2)\pi$.
If $m\geq 2$, then $|f_{0}|\geq 2m$ gives
$w(f_{0})=\Bigl(1-\frac{2}{|f_{0}|}\Bigr)\pi\ \geq\ \Bigl(1-\frac{1}{m}\Bigr)\pi,$
so
$(|\Gamma|-1)\pi+(m+1)w(f_{0})\ \geq\ (m+|\Gamma|-2)\pi+\Bigl(1-\frac{1}{m}\Bigr)\pi\ >\ (|C|-2)\pi,$
contradicting (3.2).
If $m=1$, then $|f_{0}|\geq 3$ gives $w(f_{0})\geq\pi/3$, hence
$(|\Gamma|-1)\pi+(m+1)w(f_{0})\ \geq\ (|\Gamma|-1)\pi+2\cdot\frac{\pi}{3}\ >\ (|\Gamma|-1)\pi\ =\ (|C|-2)\pi,$
again contradicting (3.2).
This contradiction proves (1a).
Proof of (1b). Let
$v:=\mathrm{first}(\partial_{R}R_{i_{s}},l_{uw}(\zeta_{i_{s}},\zeta_{i_{s+1}}]).$
Assume for contradiction that the edge along $\partial_{R}R_{i_{s}}$ immediately preceding $v$ is on the left of $l_{uw}(\zeta_{i_{s}},\zeta_{i_{s+1}}]$.
Let $\langle\zeta_{i_{s}},x_{1}\rangle$ be the distinguished incident edge on the left of $l_{uw}$ used to define $R_{i_{s}}$. Write the subpath of $l_{uw}$ from $\zeta_{i_{s}}$ to $v$ as
$\zeta_{i_{s}}(=z_{c}),z_{c+1},\dots,z_{b}(=v).$
Let $g\in\{c,\dots,b-1\}$ be maximal such that the distinguished left edge at $z_{g}$ is still $\langle z_{g},x_{1}\rangle$ (equivalently, $R$ is the last nonempty chandelier before $y$ whose distinguished neighbor off $l_{uw}$ is $x_{1}$). By construction of $\mathcal{R}_{uw}$, every vertex $z_{g+1},\dots,z_{b-1}$ is incident to no edge on the left of $l_{uw}$. Hence the whole segment
$P:=l_{uw}[z_{g},v]$
is incident (on its left side) to a single face $f_{0}$, and $y_{1}\in\partial f_{0}$. Set $m:=|E(P)|=b-g\geq 0$. Then always
(5.1) $|f_{0}|\geq\max\{m+2,2m,3\},$
Step 1: build a simple cycle. Let $\Gamma$ be the subpath of the boundary ray $\partial_{L}R_{i_{s}}$ joining $x_{1}$ to $v$ . By the choice of $v$ and the fact that boundary rays are self-avoiding, the union
$C:=P\cup\Gamma\cup\langle\zeta_{i_{s}},x_{1}\rangle$
is a simple cycle. Let $F_{C}$ be the set of faces in the bounded component of $\mathbb{R}^{2}\setminus C$.
Step 2: a unique interior face along $P$. By the construction of the points $\zeta_{i_{s}}$ on $l_{uw}$, every internal vertex of $P$ (i.e. of $P\setminus\{z_{g},v\}$) is incident to exactly one face on the left of $l_{uw}$. Since the interior of $C$ lies on the left side of $P$, it follows that the interior of $C$ is incident to every edge of $P$ through a single face, denoted $f_{0}\in F_{C}$, and $\partial f_{0}$ contains $P$ as a subwalk.
Step 3: Gauss–Bonnet contradiction. Apply Lemma 3.2 (3.2) to the simple cycle $C$:
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}\frac{|f|-2}{|f|}\pi\ \leq\ (|C|-2)\pi.$
Let $w(f):=\frac{|f|-2}{|f|}\pi$.
(i) Contribution from $P$. The face $f_{0}$ is incident to every vertex of $P$, hence for each $v\in P\cap V$,
$\sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ w(f_{0}).$
Therefore the total contribution from vertices of $P$ is at least $(m+1)w(f_{0})$.
(ii) Contribution from $\Gamma\setminus\{x_{1},v\}$. Each internal vertex of $\Gamma$ lies on a boundary ray of a chandelier, so by the defining turning rule of chandeliers the interior of $C$ occupies at least three face-sectors around such a vertex. All those faces belong to $F_{C}$ and each has degree at least $3$, hence each such vertex contributes at least $3\cdot(\pi/3)=\pi$. There are $|\Gamma|-1$ internal vertices on $\Gamma$, so their total contribution is at least $(|\Gamma|-1)\pi$.
(iii) Contribution from $x_{1}$. The vertex $x_{1}$ has degree at least 7; from definition 5.1 we see that $x_{1}$ has a unique neighbor $x_{2}$ in $R_{i_{s}}$ other than $\zeta_{i_{s}}$, and moving along $\zeta_{i_{s}}\rightarrow x_{1}\rightarrow x_{2}$, exactly 3 faces are on the left at $x_{1}$. Moreover by Lemma 5.2, $g\leq c+2$; by Lemma 3.8, every 3-cycle is a face; therefore moving along $z_{g}\rightarrow x_{1}\rightarrow x_{2}$, there are at least 2 faces on the right at $x_{1}$. Since each face has degree at least 3, the contribution of $x_{1}$ is at least $\frac{2\pi}{3}$.
Combining (i), (ii) and (iii) yields
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ (|\Gamma|-1)\pi+(m+1)w(f_{0})+\frac{2\pi}{3}.$
Since $|C|=m+|\Gamma|+1$, it remains to compare the right-hand side with $(|C|-2)\pi$.
If $m\geq 2$, then $|f_{0}|\geq 2m$ gives
$w(f_{0})=\Bigl(1-\frac{2}{|f_{0}|}\Bigr)\pi\ \geq\ \Bigl(1-\frac{1}{m}\Bigr)\pi,$
so
$(|\Gamma|-1)\pi+(m+1)w(f_{0})+\frac{2\pi}{3}\ \geq\ (m+|\Gamma|-2)\pi+\Bigl(1-\frac{1}{m}\Bigr)\pi+\frac{2\pi}{3}\ >\ (|C|-2)\pi,$
contradicting (3.2).
If $m=1$, then $|f_{0}|\geq 3$ gives $w(f_{0})\geq\pi/3$, hence
$(|\Gamma|-1)\pi+(m+1)w(f_{0})+\frac{2\pi}{3}\ \geq\ (|\Gamma|-1)\pi+2\cdot\frac{\pi}{3}+\frac{2\pi}{3}\ >\ |\Gamma|\pi\ =\ (|C|-2)\pi,$
again contradicting (3.2).
If $m=0$, $|C|=|\Gamma|+1$. Then $|f_{0}|\geq 3$ gives $w(f_{0})\geq\pi/3$, hence
$(|\Gamma|-1)\pi+(m+1)w(f_{0})+\frac{2\pi}{3}\ \geq\ |\Gamma|\pi\ >\ (|C|-2)\pi,$
again contradicting (3.2).
This contradiction proves (1b).
The remaining statements follow by the same cycle construction applied to the corresponding geodesic subsegment, with the extremal intersection chosen in the relevant direction, and with left/right interchanged for the anti-chandeliers. ∎
Lemma 5.7.
Suppose Setup 5.4 holds. Then
  1. (1)
    for every $1\leq s\leq k-1$, the edge along $\partial_{R}R_{i_{s}}$ immediately preceding $\mathrm{first}(\partial_{R}R_{i_{s}},l_{uw}(\zeta_{i_{s}},\zeta_{i_{s+1}}]\cup\partial_{L}R_{i_{s+1}})$ cannot be on the left of $\partial_{L}R_{i_{s+1}}$;
  2. (2)
    for every $2\leq s\leq k$, the edge along $\partial_{L}R_{i_{s}}$ immediately preceding $\mathrm{first}(\partial_{L}R_{i_{s}},l_{uw}[\zeta_{i_{s-1}},\zeta_{i_{s}})\cup\partial_{R}R_{i_{s-1}})$ cannot be on the right of $\partial_{R}R_{i_{s-1}}$;
  3. (3)
    for every $2\leq t\leq r$, the edge along $\partial_{R}U_{j_{t}}$ immediately preceding $\mathrm{first}(\partial_{R}U_{j_{t}},l_{uw}[\xi_{j_{t-1}},\xi_{j_{t}})\cup\partial_{L}U_{j_{t-1}})$ cannot be on the left of $\partial_{L}U_{j_{t-1}}$.
  4. (4)
    for every $1\leq t\leq r-1$, the edge along $\partial_{L}U_{j_{t}}$ immediately preceding $\mathrm{first}(\partial_{L}U_{j_{t}},l_{uw}(\xi_{j_{t}},\xi_{j_{t+1}})\cup\partial_{R}U_{j_{t+1}})$ cannot be on the right of $\partial_{R}U_{j_{t+1}}$.
Proof.
We prove the statement (1) here. The proof for statement (2)-(4) is identical after exchanging left/right throughout.
Step 1 (two consecutive chandeliers and the key face). Fix $l\in\{1,\dots,k-1\}$ and set
$R:=R_{i_{l}},\qquad R^{\prime}:=R_{i_{l+1}}.$
Let $x:=\operatorname{rt}(R)$ and $y:=\operatorname{rt}(R^{\prime})$ be their roots on $l_{uw}$, ordered so that the $z$-index of $x$ is strictly smaller than that of $y$. Let $\langle x,x_{1}\rangle$ and $\langle y,y_{1}\rangle$ be the distinguished incident edges on the left of $l_{uw}$ used to define $R$ and $R^{\prime}$ (so $x_{1},y_{1}\notin l_{uw}$). Since $R\neq R^{\prime}$, we have $x_{1}\neq y_{1}$.
Write the subpath of $l_{uw}$ from $x$ to $y$ as
$x(=z_{c}),z_{c+1},\dots,z_{b}(=y).$
Let $g\in\{c,\dots,b-1\}$ be maximal such that the distinguished left edge at $z_{g}$ is still $\langle z_{g},x_{1}\rangle$ (equivalently, $R$ is the last nonempty chandelier before $y$ whose distinguished neighbor off $l_{uw}$ is $x_{1}$). By construction of $\mathcal{R}_{uw}$, every vertex $z_{g+1},\dots,z_{b-1}$ is incident to no edge on the left of $l_{uw}$. Hence the whole segment
$P:=l_{uw}[z_{g},y]$
is incident (on its left side) to a single face $f_{0}$, and $y_{1}\in\partial f_{0}$. Set $m:=|E(P)|=b-g\geq 1$. Then always
(5.2) $|f_{0}|\geq m+2,$
since $\partial f_{0}$ contains the distinct vertices $z_{g},z_{g+1},\dots,z_{b}(=y),y_{1}$. Moreover, if $x_{1}\in\partial f_{0}$, then
(5.3) $|f_{0}|\geq m+3,$
since $\partial f_{0}$ then also contains $x_{1}$.
Step 2 (assume a positive boundary intersection and form a simple cycle). Let
$a:=\mathrm{first}(\partial_{R}R,l_{uw}(\zeta_{i_{s}},\zeta_{i_{s+1}}]\cup\partial_{L}R^{\prime}).$
Assume for contradiction that the edge along $\partial_{R}R$ immediately preceding $\mathrm{first}(\partial_{R}R,l_{uw}(\zeta_{i_{s}},\zeta_{i_{s+1}}]\cup\partial_{L}R^{\prime})$ is on the left of $\partial_{L}R^{\prime}$. Let $L$ be the ray of $\partial_{R}R$ starting from $x_{1}$ and going to infinity without passing through $x$, and let $L^{\prime}$ be the ray of $\partial_{L}R^{\prime}$ starting from $y_{1}$ and going to infinity without passing through $y$. Define the extended rays
$\widetilde{L}:=\langle z_{g},x_{1}\rangle\cup L,\qquad\widetilde{L}^{\prime}:=\langle y,y_{1}\rangle\cup L^{\prime}.$
Since $\widetilde{L}[z_{g},a]$ and $\widetilde{L}^{\prime}[y,a]$ do not hit $P$ except at their initial vertices $z_{g}$ and $y$, and boundary rays are self-avoiding, by the choice of $a$ the concatenation
[…]
is a simple cycle in the plane. Let $F_{C}$ be the set of (finite) faces in the bounded component enclosed by $C$.
Since $\deg(v)\geq 7$ and every finite face has degree at least $3$, we have $\kappa\leq 0$ (Lemma 3.5), hence Gauss–Bonnet yields
(5.4) $\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}\frac{|f|-2}{|f|}\pi\ \leq\ (|C|-2)\pi.$
Step 3 (lower bound contradicting (5.4)). For each finite face $f$, write $w(f):=\frac{|f|-2}{|f|}\pi$. Since every $f\in F_{C}$ has $|f|\geq 3$, we have $w(f)\geq\pi/3$.
(i) Contribution from $P$. The face $f_{0}$ lies inside $C$, hence $f_{0}\in F_{C}$, and $f_{0}$ is incident to every vertex of $P$. Therefore
(5.5) $\sum_{v\in P\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ (m+1)\,w(f_{0}).$
(ii) Contribution from the boundary arcs. Every vertex of $C\setminus P$ lies on $\widetilde{L}[z_{g},a]$ or on $\widetilde{L}^{\prime}[y,a]$. By the “extremal wedge” condition in Definition 5.1, the bounded region enclosed by $C$ occupies at least three face-sectors around each such vertex, except possibly at the first intersection point $a$ and (when present on $C$ as a distinct vertex) at $x_{1}$. Consequently, each non-exceptional vertex of $C\setminus P$ contributes at least $\pi$, while $a$ contributes at least $\pi/3$.
By assumption, $a\neq z_{g}$. We distinguish two possibilities for the first intersection point $a$.
Case 1: $a\notin\{z_{g},y_{1}\}$. Then $y_{1}$ is a distinct vertex of $C$, and at $y_{1}$ the interior of $C$ contains at least two faces of $F_{C}$ besides $f_{0}$, hence
$\sum_{f\in F_{C}:\,f\sim y_{1}}w(f)\ \geq\ w(f_{0})+\frac{2\pi}{3}.$
Case 1a: $x_{1}\in\partial f_{0}$. The contribution of $x_{1}$ is at least $w(f_{0})+\frac{\pi}{3}$. Using (5.3) and (5.5), and counting contributions as above,
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ (|C|-m-4)\pi\;+\;(m+3)w(f_{0})\;+\;\frac{4\pi}{3}.$
Subtracting $(|C|-2)\pi$ gives
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)-(|C|-2)\pi\ \geq\ (m+3)w(f_{0})-m\pi-\frac{2\pi}{3}.$
Since $|f_{0}|\geq m+3$, we have $w(f_{0})=\left(1-\frac{2}{|f_{0}|}\right)\pi\geq\left(1-\frac{2}{m+3}\right)\pi$, so
$(m+3)w(f_{0})-m\pi-\frac{2\pi}{3}\ \geq\ (m+3)\left(1-\frac{2}{m+3}\right)\pi-m\pi-\frac{2\pi}{3}\ =\ \frac{1}{3}\pi\;>\;0,$
contradicting (5.4).
Case 1b: $x_{1}\notin\partial f_{0}$. Then at $z_{g}$ the two edges $\langle z_{g},z_{g+1}\rangle$ and $\langle z_{g},x_{1}\rangle$ force an additional enclosed face of degree at least $3$, hence $z_{g}$ contributes at least $w(f_{0})+\pi/3$. Thus
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ (|C|-m-3)\pi\;+\;(m+2)w(f_{0})\;+\;\frac{4\pi}{3},$
and therefore
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)-(|C|-2)\pi\ \geq\ (m+2)w(f_{0})-m\pi+\frac{\pi}{3}.$
By (5.2), $(m+2)w(f_{0})\geq m\pi$, so the right-hand side is at least $\pi/3>0$, again contradicting (5.4).
Case 2: $a=y_{1}$. Then $y_{1}$ is the unique intersection point used to close the cycle.
Case 2a: $x_{1}\in\partial f_{0}$. The contribution of $x_{1}$ is at least $w(f_{0})+\frac{\pi}{3}$. Using (5.3) and the same counting as in Case 1a (but without a separate contribution from $a$), we get
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ (|C|-m-3)\pi+(m+3)w(f_{0})+\frac{\pi}{3}\;>\;(|C|-2)\pi,$
contradicting (5.4).
Case 2b: $x_{1}\notin\partial f_{0}$. As in Case 1b, the vertex $z_{g}$ contributes at least $w(f_{0})+\pi/3$. Hence
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)-(|C|-2)\pi\ \geq\ (m+2)w(f_{0})-m\pi+\frac{\pi}{3}\ \geq\ \frac{\pi}{3}\;>\;0,$
again contradicting (5.4).
All cases lead to a contradiction. Then (1) follows. ∎
Lemma 5.8.
Suppose Setup 5.4 holds. Let $x:=\zeta_{i_{1}}$ and $y:=\xi_{j_{1}}$. Assume without loss of generality that $x$ precedes $y$ along $l_{uw}$. Then the edge along $\partial_{L}R_{i_{1}}$ immediately preceding $\mathrm{first}(\partial_{L}R_{i_{1}},l_{uw}(x,y])$ cannot be on the right of $l_{uw}(x,y]$, and the edge along $\partial_{R}U_{j_{1}}$ immediately preceding $\mathrm{first}(\partial_{R}U_{j_{1}},l_{uw}[x,y))$ cannot be on the right of $l_{uw}[x,y))$. (The opposite order $y\prec x$ follows by reversing the orientation of $l_{uw}$ and exchanging left/right.)
Similarly, let $x^{\prime}:=\zeta_{i_{k}}$ and $y^{\prime}:=\xi_{j_{r}}$ and assume WLOG that $x^{\prime}$ precedes $y^{\prime}$. Then the edge along $\partial_{R}R_{i_{k}}$ immediately preceding $\mathrm{first}(\partial_{R}R_{i_{k}},l_{uw}(x^{\prime},y^{\prime}])$ cannot be on the left of $l_{uw}(x^{\prime},y^{\prime}]$, and the edge along $\partial_{R}U_{j_{r}}$ immediately preceding $\mathrm{first}(\partial_{R}U_{j_{r}},l_{uw}[x^{\prime},y^{\prime}))$ cannot be on the left of $l_{uw}[x^{\prime},y^{\prime})$. (The opposite order follows by symmetry as above.)
Proof.
This is proved by the same Gauss–Bonnet cycle argument as in Lemma 5.6: choose the first/last forbidden boundary–geodesic intersection (with respect to the orientation of $l_{uw}$), build a simple cycle using the corresponding geodesic subsegment of $l_{uw}$ and the relevant boundary ray, identify the unique “key” face along that geodesic subsegment using the extremality built into the choices of $i_{1},j_{1},i_{k},j_{r}$ in Setup 5.4, and then apply (3.2) to obtain a contradiction. We omit the repetitive details. ∎
Lemma 5.9.
Suppose Setup 5.4 holds. Then
  1. (1)
    Let $x:=\operatorname{rt}(R_{i_{1}})\in l_{uw}$ and $y:=\operatorname{rt}(U_{j_{1}})\in l_{uw}$. WLOG, assume the $z$-index of $x$ is less than or equal to that of $y$. Then
    1. (a)
      the edge along $\partial_{R}U_{j_{1}}$ immediately preceding $\mathrm{first}(\partial_{R}U_{j_{1}},l_{uw}[x,y)\cup\partial_{L}R_{i_{1}})$ cannot be on the left of $\partial_{L}R_{i_{1}}$; and
    2. (b)
      the edge along $\partial_{L}R_{i_{1}}$ immediately preceding $\mathrm{first}(\partial_{L}R_{i_{1}},l_{uw}(x,y]\cup\partial_{R}U_{j_{1}})$ cannot be on the right of $\partial_{L}U_{j_{1}}$; and
    If the $z$-index of $y$ is strictly smaller than that of $x$, the same conclusions hold after exchanging the roles of $x$ and $y$.
  2. (2)
    Let $x^{\prime}:=\operatorname{rt}(R_{i_{k}})\in l_{uw}$ and $y^{\prime}:=\operatorname{rt}(U_{j_{r}})\in l_{uw}$. WLOG, assume the $z$-index of $x^{\prime}$ is less than or equal to that of $y^{\prime}$. Then
    1. (a)
      the edge along $\partial_{R}R_{i_{k}}$ immediately preceding $\mathrm{first}(\partial_{R}R_{i_{k}},l_{uw}(x^{\prime},y^{\prime}]\cup\partial_{L}U_{j_{r}})$ cannot be on the left of $\partial_{L}U_{j_{r}}$.
    2. (b)
      the edge along $\partial_{L}U_{j_{r}}$ immediately preceding $\mathrm{first}(\partial_{L}U_{j_{r}},l_{uw}[x^{\prime},y^{\prime})\cup\partial_{R}R_{i_{k}})$ cannot be on the right of $\partial_{R}R_{i_{k}}$.
    If the $z$-index of $y^{\prime}$ is strictly smaller than that of $x^{\prime}$, the same conclusions hold after exchanging the roles of $x^{\prime}$ and $y^{\prime}$.
Proof.
We only prove (1a) and (2a) here; (1b) and (2b) can be proved similarly.
Proof of (1a). Write $U:=U_{j_{1}}$, $R:=R_{i_{1}}$, $x:=\operatorname{rt}(R)\in l_{uw}$ and $y:=\operatorname{rt}(U)\in l_{uw}$.
Let
$b:=\mathrm{first}(\partial_{R}U,l_{uw}[x,y)\cup\partial_{L}R)$
Assume for contradiction that the edge along $\partial_{R}U$ immediately preceding $b$ is on the left of $\partial_{L}R$. So $b\neq y$; and if $x\neq y$ then $b\neq x$ by Remark 5.5.
Let
$m:=|E(l_{uw}[x,y])|=d_{G}(x,y)\geq 0$
Let $P:=l_{uw}[x,y]$, and let $\partial_{R}R[x,b]$ (resp. $\partial_{L}U[y,b]$) denote the initial segment of $\partial_{R}R$ from $x$ to $b$ (resp. of $\partial_{L}U$ from $y$ to $b$). By the choice of $b$ as the first boundary intersection and by Lemma 5.8, Lemma 5.7(4) and Lemma 5.6 (2b), the three paths
$P,\qquad\partial_{R}R[x,b],\qquad\partial_{L}U[y,b]$
are internally vertex-disjoint. Hence their union
[…]
is a simple cycle. Let $F_{C}$ be the set of faces in the bounded component enclosed by $C$.
Key face along the geodesic segment. From the construction of $\mathcal{U}_{uw}$ no vertex of $P\setminus\{x,y\}$ can have an edge entering the interior of $C$. Therefore the interior of $C$ is incident to the whole path $P$ through a single face $f_{0}\in F_{C}$ whose boundary contains $P$. Since $P$ is a geodesic from $x$ to $y$, the complementary xxyy arc of $\partial f_{0}$ has length at least $m$, hence
(5.6) $|f_{0}|\geq\max\{2m,3\}\qquad(m\geq 0).$
Gauss–Bonnet lower bound. For each finite face $f$, write $w(f):=\frac{|f|-2}{|f|}\pi$. Since every finite face has degree at least $3$, we have $w(f)\geq\pi/3$. Each vertex of the boundary arcs $\partial_{R}R[x,b]\cup\partial_{L}U[y,b]$ other than the special vertices $x,y,b$ is incident to at least three faces of $F_{C}$ on the interior side (by the boundary-sector condition in Definition 5.1), hence contributes at least $\pi$. The intersection vertex $b$ contributes at least $\pi/3$.
If $m\geq 1$, every vertex of $P$ is incident to $f_{0}$, hence contributes at least $w(f_{0})$, and at endpoint $x$ there is in addition at least one more enclosed face of degree $\geq 3$ coming from the turn into the boundary arcs, so $x$ contributes at least $w(f_{0})+\pi/3$. Let $|C|$ be the number of vertices of the cycle. Counting contributions gives
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)$ $\geq(|C|-(m+2))\pi\;+\;(m+1)w(f_{0})\;+\;\frac{2\pi}{3}$
$=(|C|-2)\pi\;+\;\Bigl[(m+1)w(f_{0})-\left(m-\frac{2}{3}\right)\pi\Bigr].$
If $m=1$, then $|f_{0}|\geq 3$ and so $w(f_{0})\geq\pi/3$, hence the bracket equals $2w(f_{0})-\frac{\pi}{3}>0$. If $m\geq 2$, then (5.6) implies
$w(f_{0})=\left(1-\frac{2}{|f_{0}|}\right)\pi\geq\left(1-\frac{1}{m}\right)\pi,$
so
$(m+1)w(f_{0})-\left(m-\frac{2\pi}{3}\right)\pi\geq(m+1)\left(1-\frac{1}{m}\right)\pi-\left(m-\frac{2}{3}\right)\pi=\left(\frac{2}{3}-\frac{1}{m}\right)\pi>0.$
In all cases $m\geq 1$ we obtain
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)>(|C|-2)\pi,$
contradicting (3.2).
If $m=0$ (i.e. $x=y$), then $C$ is formed by the two boundary arcs from $x$ to $b$. All vertices of $C$ other than $x$ and $b$ contribute at least $\pi$ as above, while both $x$ and $b$ are incident to at least one face in $F_{C}$ of degree $\geq 3$, so each contributes at least $\pi/3$. Hence
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ (|C|-2)\pi+\frac{2\pi}{3}>(|C|-2)\pi,$
again contradicting (3.2).
Then (1a) follows.
Proof of (2a). Write $U^{\prime}:=U_{j_{r}}$, $R^{\prime}:=R_{i_{k}}$. $x^{\prime}:=\operatorname{rt}(R^{\prime})\in l_{uw}$ and $y^{\prime}:=\operatorname{rt}(U^{\prime})\in l_{uw}$. Let $\langle x^{\prime},x_{1}\rangle$ (resp. $\langle y^{\prime},y_{1}\rangle$) be the distinguished incident edges on the left (resp. right) of $l_{uw}$ used to define $R^{\prime}$ (resp. $U^{\prime}$); so $x_{1},y_{1}\notin l_{uw}$. Write the subpath of $l_{uw}$ from $x^{\prime}$ to $y^{\prime}$ as
$x^{\prime}(=z_{c}),z_{c+1},\dots,z_{b}(=y^{\prime}).$
Let $g_{1}\in\{c,\dots,b-1\}$ be maximal such that the distinguished left edge at $z_{g_{1}}$ is still $\langle z_{g_{1}},x_{1}\rangle$. Without loss of generality, assume $g_{1}\leq b$. By construction of $\mathcal{R}_{uw}$, every vertex $z_{g_{1}+1},\dots,z_{b-1}$ is incident to no edge on the left of $l_{uw}$. Hence the whole segment
$P:=l_{uw}[z_{g_{1}},z_{b}]$
is incident (on its left side) to a single face $f_{0}$. Set $m:=|E(P)|=b-g_{1}\geq 0$. Then always
$|f_{0}|\geq\max\{2m,3\},$
Let
$h:=\mathrm{first}(\partial_{R}R^{\prime},l_{uw}(x^{\prime},y^{\prime}]\cup\partial_{L}U^{\prime})$
Assume for contradiction that the edge along $\partial_{R}R^{\prime}$ immediately preceding $h$ is on the left of $\partial_{L}U^{\prime}$. So $h\neq x^{\prime}$; and if $x^{\prime}\neq y^{\prime}$ then $h\neq y^{\prime}$ by Remark 5.5.
By the choice of $h$ and by Lemma 5.8, Lemma 5.7(3) and Lemma 5.6 (2a), the union
[…]
is a simple cycle. Let $F_{C}$ be the set of faces in the bounded component enclosed by $C$.
Gauss–Bonnet lower bound. For each finite face $f$, write $w(f):=\frac{|f|-2}{|f|}\pi$. Since every finite face has degree at least $3$, we have $w(f)\geq\pi/3$. Each vertex of the boundary arcs $\partial_{R}R[x_{1},h]\cup\partial_{L}U[y^{\prime},h]$ other than the special vertices $x_{1},y^{\prime},h$ is incident to at least three faces of $F_{C}$ on the interior side (by the boundary-sector condition in Definition 5.1), hence contributes at least $\pi$. The intersection vertex $h$ contributes at least $\pi/3$.
If $m\geq 1$, every vertex of $P$ is incident to $f_{0}$, hence contributes at least $w(f_{0})$. The contribution at $y^{\prime}$ is at least $w(f_{0})+\frac{5\pi}{3}$
If $x_{1}\notin\partial f_{0}$ at endpoint $x$ there is in addition at least one more enclosed face of degree $\geq 3$ coming from the turn into the boundary arcs, so $x$ contributes at least $w(f_{0})+\pi/3$. From the Proof of Lemma 5.6(1b) Step 4(iii), we see the contribution at $x_{1}$ is at least $\frac{2\pi}{3}$. Let $|C|$ be the number of vertices of the cycle. Counting contributions gives
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)$ $\geq(|C|-(m+3))\pi\;+\;(m+1)w(f_{0})\;+\;3\pi$
$=(|C|-2)\pi\;+\;\Bigl[(m+1)w(f_{0})-\left(m-2\right)\pi\Bigr].$
If $_{1}\in\partial f_{0}, xx$ contributes at least $w(f_{0})$; the contribution at $x_{1}$ is at least $\frac{\pi}{3}+w(f_{0})$. Then
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)$ $\geq(|C|-(m+3))\pi\;+\;(m+2)w(f_{0})\;+\;\frac{7\pi}{3}$
$=(|C|-2)\pi\;+\;\Bigl[(m+2)w(f_{0})-\left(m-\frac{4}{3}\right)\pi\Bigr].$
If $m=1$, then $|f_{0}|\geq 3$ and so $w(f_{0})\geq\pi/3$, hence the bracket is positive. If $m\geq 2$, then (5.6) implies
$w(f_{0})=\left(1-\frac{2}{|f_{0}|}\right)\pi\geq\left(1-\frac{1}{m}\right)\pi,$
so
$(m+1)w(f_{0})-\left(m-2\right)\pi\geq(m+1)\left(1-\frac{1}{m}\right)\pi-\left(m-2\right)\pi=\left(2-\frac{1}{m}\right)\pi>0.$
and
$(m+2)w(f_{0})-\left(m-\frac{4}{3}\right)\pi\geq(m+2)\left(1-\frac{1}{m}\right)\pi-\left(m-\frac{4}{3}\right)\pi=\left(\frac{7}{3}-\frac{2}{m}\right)\pi>0.$
In all cases $m\geq 1$ we obtain
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)>(|C|-2)\pi,$
contradicting (3.2).
If $m=0$ (i.e. $x=y$), then $C$
$C=\partial_{R}R[x_{1},h]\ \cup\ \partial_{L}U[y^{\prime},h]\cup\langle z_{g_{1}},x_{1}\rangle$
All vertices of $C$ other than $z_{g_{1}}(=y^{\prime}),x_{1},h$ contribute at least $\pi$ as above, while both $z_{g_{1}}=y^{\prime}$ contributes at least $\frac{5\pi}{3}$, $h$ contributes at leastt $\pi/3$, and $x_{1}$ contributes at least $\frac{2\pi}{3}$ Hence
$\sum_{v\in C\cap V}\ \sum_{f\in F_{C}:\,f\sim v}w(f)\ \geq\ (|C|-3)\pi+\frac{8\pi}{3}>(|C|-2)\pi,$
again contradicting (3.2).
Then (2a) follows.
Recall that $G_{*}$ is the matching graph of $G$ as defined in Definition 2.1.
Lemma 5.10.
Let $G=(V,E)$ be a graph satisfying the assumptions of Definition 5.1. Then for each $p<1-p_{c}^{site}(T)$ (Recall that $1-{p_{c}^{site}}(T)>\frac{1}{2}$), there exists $c_{p}>0$, such that for any $u,v\in V$,
[…]
Moreover,
[…]
where $\partial_{V}^{*}u$ consists of all the vertices adjacent to $u$ in $G_{*}$.
Proof.
Let $l_{uv}$ be the shortest path of $G$ joining $u$ and $v$, with $u=z_{0}$ and $v=z_{n}$. Construct a sequence of chandeliers and empty sets $R_{1},R_{2},\ldots$ on the left of $l_{uv}$ and a sequence of anti-chandeliers and empty sets $U_{1},U_{2},\ldots$ on the right of $l_{uv}$.
Then one can find an alternating sequence
$R_{i_{1}}(=R_{1}),U_{j_{1}},R_{i_{2}},U_{j_{2}},\ldots$
of chandeliers, anti-chandeliers and empty sets whose roots have increasing $z$-indices, such that
  1. (A)
    The distance between the root $R_{i_{l}}$ and the root of $U_{j_{l}}$ in $G$ is at most 3; and
  2. (B)
    the distance between the root of $U_{j_{l}}$ and the root of $R_{i_{l+1}}$ in $G_{*}$ is at most 6 and at least 1;
  3. (C)
    Distinct items in the sequence are pairwise disjoint.
More precisely the sequence can be constructed as follows:
  1. (a)
    Assume $R_{i_{l}}$ is rooted at $\zeta_{i_{l}}$ and $U_{j_{l}}$ is rooted at $\xi_{j_{l}}$ along $l_{u,v}$. Then the $z$-index of $\zeta_{i_{l}}$ is less than or equal to the $z$-index of $\xi_{j_{l}}$. Moreover, there are neither elements in $\mathcal{R}_{uv}\setminus\{R_{i_{l}}\}$ nor elements in $\mathcal{U}_{uv}\setminus\{U_{j_{l}}\}$ with root along of $l_{uv}[\zeta_{i_{l}},\xi_{j_{l}}]$;
  2. (b)
    Let $R_{i_{1}},U_{j_{1}}$ be the pair satisfying condition (a) whose $\zeta_{i_{1}}$ (the root of $R_{i_{1}}$) has the minimal $z$-index; for each $l\geq 1$, let $R_{i_{l+1}},U_{j_{l+1}}$ be the pair satisfying condition (a) whose $\zeta_{i_{l+1}}$ has the minimal $z$-index strictly greater than the $z$-index of $\xi_{j_{l}}$.
Then (A) follows from Lemma 5.2, and the fact that the graph $G$ has minimal vertex degree at least 7; hence at each interior point $w$ along $l_{uv}$; either $w$ has at least two incident faces on the left of $l_{u,v}$, or $w$ has at least two incident faces on the right of $l_{uv}$.
To see why (B) is true, from (A) and (a) we see that the $z$-index of $\zeta_{i_{l}}$ is less than or equal to that of $\xi_{j_{l}}$, and moreover, $\zeta_{i_{l}}$ has maximal $z$-index among all the roots of chandeliers in $\mathcal{R}_{uv}$ whose $z$-indices do not exceed that of $\xi_{j_{l}}$; similarly, $\xi_{j_{l}}$ has minimal $z$-index among all the roots of anti-chandeliers in $\mathcal{U}_{uv}$ whose $z$-indices is greater than or equal to that of $\zeta_{i_{l}}$. By Lemma 5.2, starting from $\zeta_{i_{l}}$, moving along $l_{uv}$ to the right by at most 3 faces (with $G_{*}$ distance at most 3), there exists a chandelier $R^{\prime}$ (with root $\zeta^{\prime}$) in $\mathcal{R}_{uv}\setminus\{R_{i_{l}}\}$. Then the $z$-index of $\xi_{j_{l}}$ is strictly less than that of $\zeta^{\prime}$. Find $U^{\prime}$ in $\mathcal{U}_{uv}$ whose root $\xi^{\prime}$ has maximal $z$-index among all the roots of anti-chandeliers in $\mathcal{U}_{uv}$ whose roots have $z$-indices strictly less than that of $\zeta^{\prime}$. Again by Lemma 5.2, starting from $\xi^{\prime}$, moving along $l_{uv}$ to the right by at most 3 faces (with $G_{*}$ distance at most 3), there exists an anti-chandelier $U^{\prime\prime}$ (with root $\xi^{\prime\prime}$) in $\mathcal{U}_{uv}\setminus\{U_{i_{l}}\}$; from the choice of $U^{\prime}$ and $U^{\prime\prime}$ we see that the $z$-index of $\xi^{\prime\prime}$ is greater than or equal to that of $\zeta^{\prime}$. It follows that there exists a path from $\xi_{j_{l}}$ to $\zeta^{\prime}$ in $G_{*}$ with length at most 3, and a path form $\zeta^{\prime}$ to $\xi^{\prime\prime}$ in $G_{*}$ with length at most $3$, and the $z$-index of $\zeta_{i_{l+1}}$ is less than or equal to that of $\xi^{\prime\prime}$. Then
$d_{G_{*}}(\xi_{j_{l}},\zeta_{i_{l+1}})\leq 6$
(C) follows from Lemmas 5.7 and 5.9. See Figure 5.2.
Let $\{Q_{r}:=R_{i_{r}}\cup U_{j_{r}}\}_{r}$; then we can find at least $k:=\left\lfloor\frac{d_{G_{*}}(u,v)}{9}\right\rfloor-1$ $Q_{r}$’s along $l_{u,v}$. If $p<1-{p_{c}^{site}}(T)$, we have $1-p>{p_{c}^{site}}(T)$. Let $x_{r}$ be the root of $R_{i_{r}}$ and $y_{r}$ be the root of $U_{j_{r}}$. Then there is a positive probability $\theta_{p}>0$ such that all the following occur
  1. (1)
    $x_{r}$ is in an infinite 0-cluster of $R_{i_{r}}$; and
  2. (2)
    $y_{r}$ is in an infinite 0-cluster of $U_{j_{r}}$; and
  3. (3)
    $x_{r}$ and $y_{r}$ are joined by a 0-path along $l_{uv}$.
Let $E_{r}$ be the event that all of (1)–(3) occur. By (C) and the construction of the alternating sequence, the vertex-sets
$Q_{r}\ \cup\ \bigl(l_{uv}[x_{r},y_{r}]\cap V\bigr),\qquad r=1,2,\dots,k,$
are pairwise disjoint; hence $\{E_{r}\}_{r=1}^{k}$ are independent. Moreover, on $E_{r}$ the closed vertices contain a left–right obstruction across $l_{uv}$ (with two infinite 0-clusters attached on the left/right and a 0-path along $l_{uv}$ joining $x_{r}$ to $y_{r}$), so no open *-path can pass from $u$ to $v$. In particular,
[…]
Therefore,
[…]
where we used $k=\left\lfloor\frac{d_{G_{*}}(u,v)}{9}\right\rfloor-1\geq\frac{d_{G_{*}}(u,v)}{10}$ once $d_{G_{*}}(u,v)$ is large enough. Set $c_{p}:=-\frac{\log(1-\theta_{p})}{10}>0$. Since the desired bound is trivial for the finitely many pairs with $d_{G_{*}}(u,v)$ bounded, by decreasing $c_{p}$ if necessary we obtain (5.7) for all $u,v\in V$.
For (5.8), apply the same construction to a shortest $G_{*}$-path joining $\partial_{V}^{*}u$ to $\partial_{V}^{*}v$. This yields a constant $\varphi_{p}>0$ and at least $\frac{d_{G_{*}}(u,v)}{9}-3$ disjoint blocks producing separating events, so
[…]
after possibly decreasing $c_{p}$ once more. This proves (5.8). ∎
Remark 5.11.
The proof of Lemma 5.10 shows that when $G$ is a graph satisfying the assumptions of Definition 5.1,
$\frac{1}{2}<1-p_{c}^{\mathrm{site}}(T)\leq p_{\mathrm{exp}}(G_{*})\leq p_{\mathrm{conn}}(G_{*}),$
where
$p_{\exp}(G):=\sup\Bigl\{p\in(0,1):\ \exists\,C,\gamma\in(0,\infty)\ \text{s.t.}\ \mathbb{P}_{p}(x\leftrightarrow y)\leq Ce^{-\gamma d_{G}(x,y)},\ \forall x,y\in V\Bigr\},$
and $p_{\mathrm{conn}}(\cdot)$ is defined in (1.3).
Lemma 5.12.
Let $G=(V,E)$ be a graph satisfying the assumptions of Definition 5.1. Then for any $n\in{\mathbb{N}}$, there exists $M_{n}\in{\mathbb{N}}$, such that for any $v,w\in V$ with $d_{G}(v,w)\geq M_{n}$, there exist three trees $T_{a},T_{b},T_{c}$ rooted at $w$, each isomorphic to the tree $T$ constructed in the proof of Proposition 3.10, such that all the following conditions hold:
  • $T_{a}$ (resp. $T_{b}$) has boundaries $l_{w,1}$ and $l_{w,2}$ (resp. $l_{w,3}$ and $l_{w,4}$), where each $l_{w,i}$ is a singly infinite path starting at $w$;
  • $l_{w,2}$ and $l_{w,3}$ share an edge $\langle w,x\rangle$ for some $x\in V$;
  • the (open) region $R_{w}\subset{\mathbb{R}}^{2}$ including $T_{a}\cup T_{b}]\setminus\{l_{w,1}\cup l_{w,4}\}$ and bounded by $l_{w,1}$ and $l_{w,4}$ satisfies
    $B(v,n)\subset R_{w},$
    where
    $B(v,n):=\{u\in V:\ d_{G}(v,u)\leq n\};$
  • $T_{c}\cap R_{w}=\emptyset$.
Proof.
We first construct $T_{a},T_{b},T_{c}$.
Let $T_{1},T_{2},\ldots,T_{\deg(w)}$ be the $\deg(w)$ trees rooted at $w$, each of which is isomorphic to $T$, arranged in cyclic order around $w$. For $1\leq i\leq\deg(w)$, let $S_{i}\subset{\mathbb{R}}^{2}$ be the closed region bounded by the two boundary rays of $T_{i}$ and containing $T_{i}$. Note that ${\mathbb{R}}^{2}=\bigcup_{1\leq i\leq\deg(w)}S_{i}$, and $S_{i}\cap S_{j}=\{w\}$ if $j\notin\{i-1,i+1\}$ (with indices understood cyclically). Hence $v$ belongs to at most two of the $S_{i}$’s.
The following cases might occur.
  • $v\in S_{i}\cap S_{i+1}$. In this case, let $T_{a}:=T_{i}$, $T_{b}:=T_{i+1}$, and $T_{c}:=T_{i+3}$ (indices taken cyclically).
  • $v\in S_{i}$ and $v\notin S_{j}$ for all $j\neq i$. Let $l_{1},l_{2}$ be the two boundary rays of $T_{i}$, each a singly infinite path starting at $w$, and $l_{1}\cap l_{2}=\{w\}$. Without loss of generality, assume
    $d_{\,G\setminus B\!\left(w,\left\lfloor\frac{d_{G}(v,w)}{2}\right\rfloor\right)}(v,l_{1})\leq d_{\,G\setminus B\!\left(w,\left\lfloor\frac{d_{G}(v,w)}{2}\right\rfloor\right)}(v,l_{2}).$
    The following cases might occur.
    • the edge $\langle w,t\rangle$ along $l_{1}$ is also on the boundary of $T_{i-1}$. In this case, let $T_{a}:=T_{i-1}$, $T_{b}:=T_{i}$, and $T_{c}:=T_{i+2}$.
    • the edge $\langle w,t\rangle$ along $l_{1}$ is also on the boundary of $T_{i+1}$. In this case, let $T_{a}:=T_{i}$, $T_{b}:=T_{i+1}$, and $T_{c}:=T_{i+3}$.
Define $R_{w}$ to be the (open) region bounded by the outer boundary rays $l_{w,1}$ and $l_{w,4}$ coming from the chosen pair $(T_{a},T_{b})$, and containing $T_{a}\cup T_{b}]\setminus\{l_{w,1}\cup l_{w,4}\}$.
Now we show that
(5.11) $\lim_{d_{G}(v,w)\to\infty}d_{G}\bigl(v,\partial\overline{R_{w}}\bigr)=\infty,$
and the convergence is uniform in $w$. It is straightforward that the lemma follows from (5.11) (take $M_{n}$ so that $d_{G}(v,\partial\overline{R_{w}})>n$ whenever $d_{G}(v,w)\geq M_{n}$).
For each $k\geq 1$, let
$\partial B(w,k):=\{u\in V:\ d_{G}(w,u)=k\}.$
Without loss of generality, assume
$R_{w}=\bigl[S_{i}\cup S_{i+1}\bigr]^{\circ},$
(the case $R_{w}=[S_{i-1}\cup S_{i}]^{\circ}$ is analogous). Set
(5.12) $D_{1}(k)$ $:=d_{\,([S_{i}\cap G]\setminus B(w,k))}\bigl(l_{w,1},l_{w,2}\bigr)\ \geq\ 2^{k},$
(5.13) $D_{2}(k)$ $:=d_{\,([S_{i+1}\cap G]\setminus B(w,k))}\bigl(l_{w,3},l_{w,4}\bigr)\ \geq\ 2^{k},$
(5.14) $D_{3}(k)$ $:=d_{\,(((R_{w}\setminus S_{i})\cap G)\setminus B(w,k))}\bigl(l_{w,2},l_{w,4}\bigr)\ \geq\ 2^{k-1},$
(5.15) $D_{4}(k)$ $:=d_{\,(((R_{w}\setminus S_{i+1})\cap G)\setminus B(w,k))}\bigl(l_{w,1},l_{w,3}\bigr)\ \geq\ 2^{k-1}.$
All four inequalities can be proved similarly. For example, to see why (5.15) holds, let $w_{0}$ be the neighbor of $w$ on the boundary of $S_{i}$ that does not lie in $S_{i+1}$. Then there is a tree $T_{w_{0}}$ rooted at $w_{0}$ isomorphic to $T_{i}$ with one boundary ray given by $l_{w,1}\setminus\langle w,w_{0}\rangle$ and satisfying
$T_{w_{0}}\subset\bigl(\overline{R_{w}}\setminus S_{i+1}\bigr)\cap G.$
All vertices of $T_{i}$ whose distance (in $T_{i}$) to $w_{0}$ is at most k1k-1 are removed in $\bigl[((\overline{R_{w}}\setminus S_{i+1})\cap G)\setminus B(w,k)\bigr]$. Therefore any path joining $l_{w,1}$ and $l_{w,3}$ inside $\bigl[((\overline{R_{w}}\setminus S_{i+1})\cap G)\setminus B(w,k)\bigr]$ must pass through at least $2^{k-1}$ vertices, and (5.15) follows.
For each $k\geq 1$, define
[…]
and set
$q\bigl(v,\partial\overline{R_{w}},k\bigr):=\min\bigl\{|l|:\ l\in\mathcal{L}(v,\partial\overline{R_{w}},k)\bigr\},$
with the convention $\min\emptyset=\infty$. Then
$d_{G}\bigl(v,\partial\overline{R_{w}}\bigr)=\min_{1\leq k\leq d_{G}(v,w)}q\bigl(v,\partial\overline{R_{w}},k\bigr).$
If $1\leq k\leq\frac{d_{G}(v,w)}{2}$, then for any such SAW $l$ we must have $|l|\geq d_{G}(v,w)-k\geq\frac{d_{G}(v,w)}{2}$, hence
$q\bigl(v,\partial\overline{R_{w}},k\bigr)\geq\frac{d_{G}(v,w)}{2}.$
If $\frac{d_{G}(v,w)}{2}\leq k\leq d_{G}(v,w)$, then
$q\bigl(v,\partial\overline{R_{w}},k\bigr)$ $\geq d_{\,G\setminus B(w,k)}\bigl(v,\partial\overline{R_{w}}\bigr)$
$\geq d_{\,G\setminus B\left(w,\left\lfloor\frac{d_{G}(v,w)}{2}\right\rfloor\right)}\bigl(v,\partial\overline{R_{w}}\bigr)$
$\geq\frac{1}{2}\min\left\{D_{1}\!\left(\left\lfloor\frac{d_{G}(v,w)}{2}\right\rfloor\right),D_{2}\!\left(\left\lfloor\frac{d_{G}(v,w)}{2}\right\rfloor\right),D_{3}\!\left(\left\lfloor\frac{d_{G}(v,w)}{2}\right\rfloor\right),D_{4}\!\left(\left\lfloor\frac{d_{G}(v,w)}{2}\right\rfloor\right)\right\}$
$\geq 2^{\left\lfloor\frac{d_{G}(v,w)}{2}\right\rfloor-2}.$
Combining the two cases yields (5.11) (uniformly in $w$). In particular, choosing $M_{n}:=2n$ suffices, and the lemma follows. ∎
Theorem 5.13.
Let $G=(V,E)$ be a graph satisfying the assumptions of Definition 5.1. The following statements hold.
  1. (1)
    For each $p\in\bigl({p_{c}^{site}}(G),\,1-{p_{c}^{site}}(T)\bigr)$, there are a.s. infinitely many infinite $1$-clusters.
  2. (2)
    If $G$ has finitely many ends, for each $p\in\bigl[\,1-p_{c}(T),\,1-{p_{c}^{site}}(G_{*})\,\bigr)$, a.s. the total number of ends of all infinite $1$-clusters is infinite, i.e.
    (5.16) $\sum_{C\in\mathcal{C}_{\infty}(\omega)}\mathrm{Ends}(C)=\infty,$
    where $\mathcal{C}_{\infty}(\omega)$ denotes the collection of infinite $1$-clusters in the percolation configuration $\omega$, and $\mathrm{Ends}(C)$ is the number of ends of the (connected) graph induced by $C$.
  3. (3)
[…]
Proof.
When $p\in\bigl({p_{c}^{site}}(G),\,1-p_{c}(T)\bigr)$, the existence of infinitely many infinite $1$-clusters follows from Lemmas 1.4 and 5.10.
Now consider the case when $p\in\bigl[\,1-p_{c}(T),\,1-{p_{c}^{site}}(G_{*})\,\bigr)$. In this case when $G$ satisfies the assumptions of Definition 5.1, we have
$1-p\in\bigl({p_{c}^{site}}(G_{*}),\,p_{c}(T)\bigr]\subseteq\bigl({p_{c}^{site}}(G_{*}),\,\tfrac{1}{2}\bigr]\subseteq\bigl({p_{c}^{site}}(G_{*}),\,p_{conn}(G_{*})\bigr),$
where the last inclusion follows from Remark 5.11. Then by Lemma 1.4, a.s. there are infinitely many infinite 0-*-clusters.
Assume the graph $G$ has finitely many ends. In this case, since when $p\in\bigl[\,1-p_{c}(T),\,1-{p_{c}^{site}}(G_{*})\,\bigr)$ infinite 0-*-clusters have infinitely many ends, we infer that infinite $1$-clusters have infinitely many ends to separate the infinitely many ends of infinite 0-*-clusters. This completes the proof of Part (2).
Now we prove Part (3). Assume that $G$ has infinitely many ends. Fix a vertex $v_{0}\in V$. Recall that $B(v_{0},n)$ is the ball consisting of all vertices within graph distance $n$ of $v_{0}$ in $G$. Let $G\setminus B(v_{0},n)$ be the subgraph of $G$ obtained by removing all vertices in $B(v_{0},n)$ and all edges incident to at least one vertex in $B(v_{0},n)$. Then the number of infinite components of $G\setminus B(v_{0},n)$ tends to infinity as $n\to\infty$.
We first record the following claim.
Claim 5.14.
Let $H$ be an arbitrary infinite component of $G\setminus B(v_{0},n)$. Then we can find a tree $T_{H}$ embedded in $H$ which is isomorphic to the tree $T$ constructed in the proof of Proposition 3.10.
Proof of Claim 5.14.
Let $u$ be a vertex in $H$ adjacent to a vertex in $B(v_{0},n)$. Then $d_{G}(u,v_{0})=n+1$. Since $H$ is infinite and connected, we can find a directed singly infinite path in $H$ starting at $u$, denoted by
$z_{0}(:=u),z_{1},z_{2},\ldots.$
Since $G$ is locally finite, we have $\lim_{m\to\infty}d_{G}(v_{0},z_{m})=\infty$. Choose $k$ so that $d_{G}(z_{k},v_{0})\geq M_{n}$, where $M_{n}$ is given by Lemma 5.12. By Lemma 5.12, there exists a tree $T_{H}$ isomorphic to the one constructed in the proof of Proposition 3.10, such that $T_{H}\cap B(v_{0},n)=\emptyset$. Since $T_{H}$ is connected and contains $z_{k}\in H$, we have $T_{H}\subseteq H$. ∎
Fix $n\geq 1$, and let $\mathcal{H}_{n}$ denote the set of infinite components of $G\setminus B(v_{0},n)$. For each $H\in\mathcal{H}_{n}$, pick one such embedded copy $T_{H}\subseteq H$ given by Claim 5.14. Since $p\in[\,1-p_{c}(T),1\,]$ and $1-p_{c}(T)>\frac{1}{2}>{p_{c}^{site}}(T)$, we have $p>{p_{c}^{site}}(T)$. Therefore, for each fixed $H\in\mathcal{H}_{n}$, the i.i.d. Bernoulli($p$) site percolation restricted to $T_{H}$ a.s. contains an infinite $1$-cluster (in $T_{H}$, hence also in $G$). In particular, a.s. for every $H\in\mathcal{H}_{n}$ there exists an infinite $1$-cluster of $G$ whose intersection with $H$ contains an infinite connected subgraph.
Consequently, a.s. the subgraph induced by the union of all infinite $1$-clusters has, after deleting $B(v_{0},n)$, at least $|\mathcal{H}_{n}|$ infinite connected components (at least one inside each $H\in\mathcal{H}_{n}$). On the other hand, if we denote by $\mathcal{C}_{\infty}(\omega)$ the set of infinite $1$-clusters in $\omega$, then for every finite set $K$ and every $C\in\mathcal{C}_{\infty}(\omega)$, the graph $C\setminus K$ has at most $\mathrm{Ends}(C)$ infinite components by the definition of ends. Hence
$\{\text{infinite components of }(\cup_{C\in\mathcal{C}_{\infty}(\omega)}C)\setminus K\}\leq\sum_{C\in\mathcal{C}_{\infty}(\omega)}\mathrm{Ends}(C).$
Applying this with $K=B(v_{0},n)$ and using the lower bound $\geq|\mathcal{H}_{n}|$, we get
$\sum_{C\in\mathcal{C}_{\infty}(\omega)}\mathrm{Ends}(C)\ \geq\ |\mathcal{H}_{n}|.$
Finally, since $G$ has infinitely many ends, $|\mathcal{H}_{n}|\to\infty$ as $n\to\infty$, and thus
$\sum_{C\in\mathcal{C}_{\infty}(\omega)}\mathrm{Ends}(C)=\infty\qquad\text{a.s.}$
This proves Part (3) and completes the proof of the theorem. ∎
6. From infinitely many ends to infinitely many infinite clusters
From Theorem 5.13, we know that when $\frac{1}{2}<p<1-p_{c}^{site}(G_{*})$, $\mathbb{P}_{p}$-a.s. the total number of ends of infinite 1-clusters is infinite. In this section we show that, under a mild hypothesis on long non-self-touching polygons in $G_{*}$ (Assumption 6.3), there exists $\delta>0$ such that for all
$1-p_{c}^{site}(G_{*})-\delta<p<1-p_{c}^{site}(G_{*}),$
$\mathbb{P}_{p}$-a.s. there are infinitely many infinite $1$-clusters. Note that by Lemmas 1.4 and 5.10, when $p\in(\tfrac{1}{2},1-p_{c}^{site}(G_{*}))$ there are $\mathbb{P}_{p}$-a.s. infinitely many infinite 0-*-clusters.
Definition 6.1.
Let $G=(V,E)$ be an infinite, connected, planar, simple graph properly embedded into ${\mathbb{R}}^{2}$ with minimal vertex degree at least $7$. Fix once and for all a total order on $V$ (e.g. by enumerating vertices).
Let $\omega\in\{0,1\}^{V}$ be a site-percolation configuration on $G$ and assume that $\xi$ is a $1$-ended infinite 0-*-cluster in $\omega$.
Define $\mathcal{I}_{\xi}$ to be the collection of doubly infinite walks
(6.1) $I_{\xi}:=\ldots,w_{-n},\ldots,w_{-1},w_{0},w_{1},\ldots,w_{n},\ldots$
in the matching graph $G_{*}$ satisfying:
  • $w_{i}\in\xi$ for all $i\in{\mathbb{Z}}$;
  • $w_{i}$ and $w_{i+1}$ are *-adjacent for all $i\in{\mathbb{Z}}$;
  • orient $I_{\xi}$ by increasing indices. Then every vertex that is *-adjacent to $I_{\xi}$ and lies on the right side of the oriented walk (with respect to the fixed planar embedding) has state $1$ in $\omega$;
  • the total number of distinct vertices visited by $I_{\xi}$ is infinite.
Definition 6.2.
Let $H$ be a graph and let $v_{0},v_{1},\ldots,v_{n}$ be a walk in $H$. A touching pair of this walk is a pair $(v_{i},v_{j})$ with $0\leq i<j\leq n$, $j-i\geq 2$, and $d_{H}(v_{i},v_{j})=1$.
An $n$-step non-self-touching walk on $H$ is an $n$-step self-avoiding walk with no touching pairs.
A non-self-touching polygon of length $n$ is a cycle $v_{0},v_{1},\ldots,v_{n}(=v_{0})$ such that for any $0\leq i<j\leq n$ with $n-2\geq j-i\geq 2$ we have $d_{H}(v_{i},v_{j})\geq 2$.
Assumption 6.3.
There exists $\delta>0$ such that for every $p\in\bigl(p_{c}^{site}(G_{*}),\,p_{c}^{site}(G_{*})+\delta\bigr)$ and every pair of vertices $u,v\in V$ with $d_{G_{*}}(u,v)=1$,
[…]
Lemma 6.4.
Let $G$ satisfy the assumptions of Definition 5.1 and assume Assumption 6.3 holds with some $\delta>0$. Then for every $\in\bigl(1-p_{c}^{site}(G_{*})-\delta,\ 1-p_{c}^{site}(G_{*})\bigr), p\mathbb{P}_{p}$-a.s. not every end of infinite 0-*-clusters is isolated. In particular, this implies that there exists an infinite 0-*-cluster with infinitely many ends.
Proof.
First of all, by the same arguments as on page 76 of [3], not every end of infinite 0-*-clusters is isolated implies that there exists an infinite 0-*-cluster with $2^{\aleph_{0}}$ ends; here $\aleph_{0}$ is the cardinality of any countably infinite set. Hence it suffices to show that not every end of infinite 0-*-clusters is isolated. Since this is a tail-event, the Kolmogorov 0-1 law implies that it has probability of either 0 or 1.
Fix $p\in\bigl(1-p_{c}^{site}(G_{*})-\delta,\ 1-p_{c}^{site}(G_{*})\bigr)$ and set q:=1pq:=1-p. Then $q\in(p_{c}^{site}(G_{*}),\,p_{c}^{site}(G_{*})+\delta)$. By Remark 5.11 we have $q\in(p_{c}^{site}(G_{*}),p_{conn}(G_{*}))$, hence by Lemma 1.4 there are $\mathbb{P}_{p}$-a.s. infinitely many infinite 0-*-clusters.
Assume for contradiction that a.s.,
(6.2) every end of infinite 0-*-clusters is isolated.
Let $\overline{G}=(\overline{V},\overline{E})$ be a graph constructed from $G$ as follows: $\overline{V}$ is obtained from $V$ by adding an extra vertex in each finite face, $\overline{E}$ is obtained from $E$ by adding an extra edge between each vertex in $\overline{V}\setminus{V}$ (i.e. an added vertex in a finite face $F$) and a vertex on the boundary of the face $F$. Then $\overline{G}$ is an infinite, connected, locally finite graph that can be properly embedded into ${\mathbb{R}}^{2}$. For site configuration $\omega\in\{0,1\}^{V}$, define a site configuration $\overline{\omega}\in\{0,1\}^{\overline{V}}$ by letting $\overline{\omega}(v)=\omega(v)$ if $v\in V$ and $\omega(v)=0$ if $v\in\overline{V}\setminus V$. Let $\overline{G}_{\omega}$ be the induced subgraph of $\overline{G}$ by state-0 vertices in $\overline{\omega}$. Then infinite 0-*-clusters in $\omega$ are in 1-1 correspondence with infinite components in $\overline{G}_{\omega}$. When (6.2) holds for $\omega$, every end of $\overline{G}_{\omega}$ is isolated.
Since $\overline{G}_{\omega}$ is a locally finite planar graph, it admits a Freudenthal embedding into the sphere $\mathbb{S}^{2}$ (see Section 2 of [20]; see also Section 8.6 of [8], Proposition 1.22 in [23]); and one can define a topology on the union $|\overline{G}_{\omega}|$ of $\overline{G}_{\omega}$ and its ends, such that the Fredenthal embedding of $|\overline{G}_{\omega}|$ into $\mathbb{S}^{2}$ is a homeomorphism from $|\overline{G}_{\omega}|$ onto its image. Under this embedding, each end of $|\overline{G}_{\omega}|$ corresponds an accumulation point of the embedding. Since $\mathbb{S}^{2}$ is first countable and each end of $|\overline{G}_{\omega}|$ is isolated, in $\omega$ infinite 0-*-clusters have countably many ends if (6.2) holds.
Moreover, on (6.2) by finite energy, with strictly positive probability there exists at least one $1$-ended infinite 0-*-cluster. For $v\in V$, let $\mathcal{E}_{v}$ be the event that $v$ belongs to a $1$-ended infinite 0-*-cluster. Then $\mathbb{P}_{p}(\bigcup_{v\in V}\mathcal{E}_{v})>0$, so by countability there exists $v_{0}\in V$ such that
(6.3) $\mathbb{P}_{p}(\mathcal{E}_{v_{0}})=c_{0}>0.$
On $\mathcal{E}_{v_{0}}$, write $\xi:=C_{0*}(v_{0})$; i.e. $\xi$ is the 1-ended infinite 0-*-cluster at $v_{0}$.
Step 1: A forbidden configuration forces arbitrarily long 0-*-polygons. Fix $n\geq 1$. For $I_{\xi}\in\mathcal{I}_{\xi}$, define
(6.4) $\Phi(I_{\xi,n}):=\{(u,v)\in[I_{\xi}\cap V]\times[I_{\xi}\cap V]:u=w_{i},\ i<-n,\ v=w_{j},\ j>n\}.$
Claim 6.5.
Let $n\geq 1$. On $\mathcal{E}_{v_{0}}$, let $\xi$ is the 1-ended infinite 0-*-cluster at $v_{0}$. On the event
$\mathcal{E}_{v_{0}}\ \cap\ \Bigl\{\exists I_{\xi}\in\mathcal{I}_{\xi}\text{ such that }\forall(u,v)\in\Phi(I_{\xi,n}),\ d_{G_{*}}(u,v)>2\Bigr\},$
there exist adjacent vertices $x,y\in V$ (equivalently an edge $\langle x,y\rangle\in E_{*}$) such that for every $M\geq 1$,
[…]
We postpone the proof of Claim 6.5.
Step 2: Use Assumption 6.3 to rule out the forbidden configuration. Since $q=1-p\in(p_{c}^{site}(G_{*}),p_{c}^{site}(G_{*})+\delta)$, Assumption 6.3 applies at parameter $q$. Moreover, for any adjacent $x,y$,
[…]
because under $\mathbb{P}_{p}$ the 0-vertices have the same law as the $1$-vertices under $\mathbb{P}_{1-p}=\mathbb{P}_{q}$.
By Claim 6.5 and a union bound over edges of $G_{*}$,
$\mathbb{P}_{p}\Bigl(\mathcal{E}_{v_{0}}\cap\{\exists I_{\xi}\in\mathcal{I}_{\xi}:\forall(u,v)\in\Phi(I_{\xi,n}),\ d_{G_{*}}(u,v)>2\}\Bigr)$
[…]
where the last equality is Assumption 6.3.
Since $n$ was arbitrary, we deduce that on $\mathcal{E}_{v_{0}}$, for every $n\geq 1$ and every $I_{\xi}\in\mathcal{I}_{\xi}$, there exists $(u,v)\in\Phi(I_{\xi,n})$ with $d_{G_{*}}(u,v)\leq 2$.
Step 3: Close pairs yield bounded cutsets and $p_{c}^{site}(\xi)=1$. Fix a $1$-ended infinite 0-*-cluster $\xi$ and $I_{\xi}\in\mathcal{I}_{\xi}$. From Steps 1–2, for infinitely many $n$ there exists $(u,v)\in\Phi(I_{\xi,n})$ with $d_{G_{*}}(u,v)\leq 2$. Choose such pairs $(u_{m},v_{m})$ so that all $u_{m},v_{m}$ are pairwise distinct (possible since the indices tend to $\pm\infty$).
For each $m$, let $P_{m}$ be a shortest $G_{*}$-path from $u_{m}$ to $v_{m}$ (so $|E(P_{m})|\leq 2$), chosen to lie on the right side of the oriented walk $I_{\xi}$. By the defining property of $\mathcal{I}_{\xi}$, the interior vertices of $P_{m}$ (if any) have state $1$, hence $P_{m}$ intersects $\xi$ only at its endpoints. The union of $P_{m}$ with the subwalk of $I_{\xi}$ from $u_{m}$ to $v_{m}$ contains a simple cycle $C_{m}$ in the plane, whose bounded component contains only finitely many vertices of $G$ (proper embedding). Consequently, any infinite *-path in $\xi$ from $v_{0}$ to infinity must cross $C_{m}$, and since $\xi\cap C_{m}=\{u_{m},v_{m}\}$, the set $\{u_{m},v_{m}\}$ is a vertex cutset in $\xi$ separating $v_{0}$ from infinity.
Thus $\xi$ contains infinitely many pairwise disjoint vertex cutsets of size at most $2$. It follows that $p_{c}^{site}(\xi)=1$: for any $s<1$, in Bernoulli($s$) site percolation on $\xi$, the events “the $m$-th cutset contains an open vertex” are independent and have probability at most $1-(1-s)^{2}<1$, hence the probability of crossing all cutsets is 0.
Step 4: Sprinkling and contradiction. Assume (6.2) holds ${\mathbb{P}}_{p}$ a.s.
Choose $p^{\prime}>p$ so that $q^{\prime}:=1-p^{\prime}>p_{c}^{site}(G_{*})$, and couple $\omega_{p},\omega_{p^{\prime}}$ by i.i.d. uniforms $(U_{v})_{v\in V}$. Conditional on $\omega_{p}$, passing from $p$ to $p^{\prime}$ performs independent thinning of the 0-vertices with retention probability $q^{\prime}/q<1$. If ${\mathbb{P}}_{p^{\prime}}$-a.s. there exists an infinite 0-*-cluster, then with strictly positive probability there exists a finite set of vertices $K$ such that removing $K$, there is an 1-ended infinite 0-*-cluster at level $p$ including an infinite 1-*-cluster at level $p^{\prime}$. But this has probability 0 by Step 3 and the fact that the total number of ends (all isolated) of infinite 0-*-clusters is countable.
This contradicts $q^{\prime}>p_{c}^{site}(G_{*})$. More precisely, by the definition of $p_{c}^{site}$ the Bernoulli($q^{\prime}$) percolation on $G_{*}$ has an infinite 11-*-cluster almost surely (equivalently, $\omega_{p^{\prime}}$ has an infinite 0-*-cluster almost surely). Hence (6.2) has probability 0. ∎
Proof of Claim 5.5. Fix $n\geq 1$ and work on the event
$\mathcal{E}_{v_{0}}\ \cap\ \Big\{\exists\,I_{\xi}\in\mathcal{I}_{\xi}\ \text{s.t.}\ d_{G^{*}}(u,v)>2\ \ \forall(u,v)\in\Phi(I_{\xi,n})\Big\}.$
Let $\xi=C_{0^{*}}(v_{0})$ be the (infinite) 0-*-cluster of $v_{0}$ and choose
$I_{\xi}=\big(\ldots,w_{-1},w_{0},w_{1},\ldots\big)\in\mathcal{I}_{\xi}$
satisfying $d_{G^{*}}(u,v)>2$ for all $(u,v)\in\Phi(I_{\xi,n})$.
Step 1: A shortcut path and the “shielded” subpath. Since $\xi$ is connected in $G^{*}$, there exists a 0-*-walk in $\xi$ joining $w_{-n-1}$ to $w_{n+1}$. By applying loop-erasure and then repeatedly shortcutting touching pairs (Definition 6.2) to this finite walk, we may assume it is non-self-touching; denote it by
$\theta_{n}=(z_{0},z_{1},\ldots,z_{k}),\qquad z_{0}=w_{-n-1},\ z_{k}=w_{n+1}.$
We now locate the first/last places where the tails of $I_{\xi}$ come within *-distance 11 from $\theta_{n}$.
Define
$\Gamma_{1}:=\Big\{m<-n-1:\ \exists\,i\in\{0,\ldots,k\}\text{ with }d_{G_{*}}(w_{m},z_{i})\leq 1\Big\},\qquad Q_{1}:=\begin{cases}\min\Gamma_{1},&\Gamma_{1}\neq\emptyset,\\ -n-1,&\Gamma_{1}=\emptyset.\end{cases}$
Then for every $r<Q_{1}$, the vertex $w_{r}$ is not *-adjacent to (and not equal to) any vertex of $\theta_{n}$.
Let $a$ be the maximal index such that $d_{G_{*}}(w_{Q_{1}},z_{a})\leq 1$. (If $Q_{1}=-n-1$, set $a=0$ so $z_{a}=z_{0}$.) Thus $w_{Q_{1}}$ is *-adjacent to $z_{a}$ but is not *-adjacent to any $z_{j}$ with $j>a$.
Next, restrict to the tail $\theta_{n,[a,k]}=(z_{a},z_{a+1},\ldots,z_{k})$ and define
$\Gamma_{2}:=\Big\{s>n+1:\ \exists\,j\in\{a,\ldots,k\}\text{ with }d_{G_{*}}(w_{s},z_{j})\leq 1\Big\},\qquad Q_{2}:=\begin{cases}\max\Gamma_{2},&\Gamma_{2}\neq\emptyset,\\ n+1,&\Gamma_{2}=\emptyset.\end{cases}$
Let $b$ be the minimal index in $\{a,\ldots,k\}$ such that $d_{G_{*}}(w_{Q_{2}},z_{b})\leq 1$. (If $Q_{2}=n+1$, set $b=k$ so $z_{b}=z_{k}$.)
We claim that $b\geq a+1$. Indeed, if $b=a$, then both $w_{Q_{1}}$ and $w_{Q_{2}}$ are within *-distance 11 from the same vertex $z_{a}$, hence $d_{G_{*}}(w_{Q_{1}},w_{Q_{2}})\leq 2$. But $Q_{1}\leq-n-1$ and $Q_{2}\geq n+1$ imply $(w_{Q_{1}},w_{Q_{2}})\in\Phi(I_{\xi,n})$, contradicting the standing assumption that $d_{G_{*}}(u,v)>2$ for all $(u,v)\in\Phi(I_{\xi,n})$. Therefore $b\geq a+1$, so the subpath $\theta_{n,[a,b]}$ contains at least one edge.
Step 2: Splicing and producing a bi-infinite non-self-touching walk. Form a bi-infinite 0-*-walk $W$ by following $I_{\xi}$ from -$\infty$ to $w_{Q_{1}}$, then (if needed) one *-edge from $w_{Q_{1}}$ to $z_{a}$, then $\theta_{n,[a,b]}$, then (if needed) one *-edge from $z_{b}$ to $w_{Q_{2}}$, and finally $I_{\xi}$ from $w_{Q_{2}}$ to $\infty$. All vertices of $W$ lie in $\xi$, hence are 0 in the *-graph.
Now simplify $W$ as follows: (i) erase loops (whenever the walk visits the same vertex twice, delete the enclosed cycle); (ii) iteratively shortcut touching pairs (Definition 6.2) until none remain. This yields a doubly infinite non-self-touching 0-*-walk, denote it by $\widetilde{I}_{\xi}$.
By the construction of $Q_{1},Q_{2},a,b$, no vertex of $\theta_{n,[a,b]}$ is *-adjacent to the part of $I_{\xi}$ strictly before $w_{Q_{1}}$ or strictly after $w_{Q_{2}}$, except through the two junctions at $w_{Q_{1}}$ and $w_{Q_{2}}$. Consequently, the only way the simplification could completely remove the shortcut $\theta_{n,[a,b]}$ would be to create a shortcut of length at most $2$ in $G_{*}$ between some $w_{i}$ with i<ni<-n and some $w_{j}$ with $j>n$, contradicting again $d_{G_{*}}(u,v)>2$ for all $(u,v)\in\Phi(I_{\xi,n})$. Therefore,
$\theta_{n,[a,b]}\ \subset\ \widetilde{I}_{\xi}\quad\text{as a subwalk.}$
Pick an edge $e=\langle x,y\rangle\in E_{*}\cap\theta_{n,[a,b]}$.
Step 3: From a doubly infinite walk to arbitrarily long non-self-touching polygons. For each $M\geq 1$, let $I_{\xi,e,M}$ be the finite subwalk of $\widetilde{I}_{\xi}$ obtained by taking $M$ steps in each direction starting from the edge $e$; thus $I_{\xi,e,M}$ is a non-self-touching 0-*-walk containing $e$ and having length $\geq 2M$.
Let $F_{M}$ be the finite vertex set consisting of (i) all vertices of $I_{\xi,e,M}$ and (ii) all vertices that are *-adjacent to an interior vertex of $I_{\xi,e,M}$. Since $\xi$ is 1-ended, removing the finite set $F_{M}$ leaves a unique infinite component; let $\widetilde{\xi}_{M}$ be this component.
Let $u$ and $v$ be the two endpoints of $I_{\xi,e,M}$, and let $u_{1}$ (resp. $v_{1}$) be the neighbors of $u$ (resp. $v$) along $\widetilde{I}_{\xi}$ that lie outside $I_{\xi,e,M}$. Because $\widetilde{I}_{\xi}$ is non-self-touching, the tails $\widetilde{I}_{\xi}\setminus I_{\xi,e,M}$ lie in $\widetilde{\xi}_{M}$, hence $u_{1},v_{1}\in\widetilde{\xi}_{M}$. Choose a (finite) 0-*-path $t_{M}$ in $\widetilde{\xi}_{M}$ connecting $u_{1}$ to $v_{1}$. By the definition of $F_{M}$, no vertex of $t_{M}$ is *-adjacent to any interior vertex of $I_{\xi,e,M}$.
Concatenate the walk
$u\to\cdots\to v\to v_{1}\xrightarrow{\,t_{M}\,}u_{1}\to u,$
which is a closed 0-*-walk containing the edge $e$. Finally, apply loop-erasure and touching-pair deletion to this closed walk. Since $t_{M}$ stays at *-distance at least $2$ from the interior of $I_{\xi,e,M}$, the shortcutting procedure cannot remove the portion of $I_{\xi,e,M}$ around $e$, and we obtain a non-self-touching 0-*-polygon $P_{M}$ that still contains $e$ and has length at least $2M$.
[…]
Lemma 6.6.
Let $G=(V,E)$ be an infinite, connected, planar graph properly embedded into ${\mathbb{R}}^{2}$ with minimal vertex degree at least $7$. Suppose Assumption 6.3 holds. Then there exists $\delta>0$ such that for each $\in\bigl(1-p_{c}^{site}(G_{*})-\delta,\ 1-p_{c}^{site}(G_{*})\bigr), p\mathbb{P}_{p}$-a.s. there are infinitely many infinite $1$-clusters.
Proof.
Fix $p$ in the stated interval.
Case 1: $G$ has finitely many ends. Choose a finite subgraph $K$ such that $G\setminus K$ has finitely many infinite components $H_{1},\ldots,H_{n}$ and each infinite component is $1$-ended. Then
$p_{c}^{site}(G_{*})=\min_{1\leq i\leq n}\{p_{c}^{site}([H_{i}]_{*})\}$
Let $H\in\{H_{1},\ldots,H_{n}\}$, such that $p_{c}^{site}(G_{*})=p_{c}^{site}(H_{*})$. Then $H$ is an infinite, connected, one-ended graph properly embedded into ${\mathbb{R}}^{2}$. We apply Lemma 6.4 to $H$, since $H$ is one-ended, ${\mathbb{P}}_{p}$-a.s. there are infinitely many infinite 1-clusters in $H$. Since $H$ is a component obtained by removing a finite set of vertices $K$ from $G$, we infer that ${\mathbb{P}}_{p}$-a.s. there are infinitely many infinite 1-clusters in $G$.
Case 2: $G$ has infinitely many ends. Fix $v_{0}\in V$ and set $K_{i}:=B(v_{0},i)$. Then the number $N_{i}$ of infinite components of $G\setminus K_{i}$ satisfies $N_{i}\to\infty$.
By Claim 5.14 (shrinking $\delta$ if needed so that $p>p_{c}^{site}(T)$), for every $i$ and every infinite component $H$ of $G\setminus K_{i}$, the restriction of $\omega$ to $H$ contains an infinite $1$-cluster almost surely. Let $E_{i}$ be the event that every infinite component of $G\setminus K_{i}$ contains an infinite $1$-cluster. Then
(6.5) ${\mathbb{P}}_{p}(E_{i})=1\qquad\forall i\geq 1.$
Assume for contradiction that $\omega$ has only finitely many infinite $1$-clusters. Let $\Theta$ be the union of all infinite 0-*-clusters together with all *-edges joining two vertices in the same infinite 0-*-cluster. By Lemma 6.4, ${\mathbb{R}}^{2}\setminus\Theta$ has infinitely many unbounded connected components.
Let $\mathcal{U}_{\infty}$ be the set of unbounded connected components of ${\mathbb{R}}^{2}\setminus\Theta$, and let
$\mathcal{U}_{1}:=\{U\in\mathcal{U}_{\infty}:\ U\text{ contains (equivalently, intersects) an infinite $1$-cluster}\}.$
Since we assumed that there are only finitely many infinite $1$-clusters and each infinite $1$-cluster is contained in a single component of ${\mathbb{R}}^{2}\setminus\Theta$, the set $\mathcal{U}_{1}$ is almost surely finite.
Claim 6.7.
On the event that $\mathcal{U}_{\infty}$ is infinite while $\mathcal{U}_{1}$ is finite, there exists $i$ and an infinite component $H$ of $G\setminus K_{i}$ such that $H$ contains no infinite $1$-cluster.
Assuming Claim 6.7, we get the desired contradiction: on $\bigcap_{i}E_{i}$ (which has probability $1$ by (6.5)), every infinite component of $G\setminus K_{i}$ contains an infinite $1$-cluster for every $i$, whereas Claim 6.7 produces an $i$ and an infinite component $H$ of $G\setminus K_{i}$ containing no infinite $1$-cluster. Therefore there must be infinitely many infinite $1$-clusters. ∎
Proof of Claim 6.7 Assume that $\mathcal{U}_{\infty}$ is infinite while $\mathcal{U}_{1}$ is finite. Set $U^{\star}:=\bigcup_{U\in\mathcal{U}_{1}}U$.
We first note that any infinite $1$-cluster $C$ is contained in a single component of ${\mathbb{R}}^{2}\setminus\Theta$: indeed, viewing $C$ as a connected union of embedded edges of $G$, it is disjoint from $\Theta$ (since $\Theta$ consists of 0-vertices and *-edges between them), and it cannot cross any *-edge of $\Theta$ by planarity of the embedding of $G_{*}$. Hence $C\subset U$ for some $U\in\mathcal{U}_{1}$.
Let $\xi$ be an infinite 0-*-cluster. We say that $\mathcal{U}_{1}$ is adjacent to an end $\phi$ of $\xi$, if for any finite subset of vertices $K$ of $V$, there is a vertex in $\mathcal{U}_{1}$ *-adjacent to a vertex in $\phi(K)$ (an infinite component of $G\setminus K$ assigned by the mapping $\phi$).
[…]
7. Non-self-touching polygons
In this section we verify Assumption 6.3 for properly embedded planar graphs whose every finite face is a triangle. The argument is a Peierls-type counting estimate for long non-self-touching polygons in the matching graph, combined with an isoperimetric inequality controlling the size of the interior neighborhood of a polygon. Related counting ideas for self-avoiding polygons on planar graphs appear in [24].
Definition 7.1.
Given a locally finite, simple graph $G=(V,E)$ and an edge $e=\langle x,y\rangle\in E$, let $c_{n}(e)$ be the number of length-$n$ non-self-touching polygons on $G$ containing $e$. Define the exponential growth rate
$\gamma_{G}(e):=\limsup_{n\to\infty}\bigl[c_{n}(e)\bigr]^{1/n}.$
Lemma 7.2.
Suppose that $G=(V,E)$ is an infinite, locally finite, connected graph and $e=\langle u,v\rangle\in E$. Then for any $p<1/\gamma_{G}(e)$,
[…]
Proof.
Let $\gamma:=\gamma_{G}(e)$. If $\gamma=0$, then $c_{n}(e)=0$ for all sufficiently large $n$, and (7.1) is trivial. Assume $\gamma\in(0,\infty)$ and fix $p<1/\gamma$. Choose $\varepsilon>0$ such that $p(\gamma+\varepsilon)<1$.
By the definition of $\limsup$, there exists $n_{0}\in\mathbb{N}$ such that $c_{n}(e)\leq(\gamma+\varepsilon)^{n}$ for all $n\geq n_{0}$. Hence for any $m\geq n_{0}$,
[…]
Since $p(\gamma+\varepsilon)<1$, the right-hand side tends to 0 as $m\to\infty$, which proves (7.1). ∎
Lemma 7.3.
Let $G=(V,E)$ be an infinite, connected graph properly embedded into $\mathbb{R}^{2}$. Assume the minimal vertex degree is at least $d$ and each finite face of $G$ is a triangle. Let
$r=\frac{1}{d-5}.$
Then for any $e\in E$,
$\gamma_{G_{*}}(e)\leq\frac{(1+r)^{1+r}}{r^{r}}.$
Proof.
Since every finite face of $G$ is a triangle, the matching graph adds no extra edges (Definition 2.1), hence $G_{*}=G$. It suffices to bound $\gamma_{G}(e)$.
Fix a non-self-touching polygon $P$ in $G$ and write $|P|=n$ for its length. Let $\mathrm{Int}(P)$ be the bounded component of $\mathbb{R}^{2}\setminus P$, and define
$\partial^{V,\circ}P:=\bigl\{x\in V\cap\mathrm{Int}(P):\ x\sim v\ \text{for some }v\in V(P)\bigr\}.$
(When all finite faces are triangles, this is equivalently the set of vertices in $\mathrm{Int}(P)$ that share a triangular face with a vertex of $P$.)
A linear isoperimetric estimate for triangulations with minimum degree at least $d$ (see, e.g., the argument of [24, Theorem 6]) yields
(7.2) $(d-5)\,|\partial^{V,\circ}P|\leq|P|=n.$
Now consider Bernoulli($p$) site percolation on $V$. We say that $P$ occurs if all vertices of $P$ are open and all vertices in $\partial^{V,\circ}P$ are closed. Then
$\mathbb{P}_{p}(P\ \text{occurs})=p^{|P|}(1-p)^{|\partial^{V,\circ}P|}\geq p^{n}(1-p)^{n/(d-5)}=\bigl[p(1-p)^{r}\bigr]^{n}.$
Fix an edge $e\in E$ and let $N_{n}(e)$ be the number of occurring length-$n$ non-self-touching polygons containing $e$. We claim that
(7.3) $N_{n}(e)\leq 2\qquad\text{for every percolation configuration.}$
Indeed, an occurring polygon containing $e$ has a well-defined interior side of $e$ (left or right in the embedding). We show there is at most one occurring polygon containing $e$ with a prescribed choice of the interior side, and hence at most two in total. Suppose for contradiction that $P\neq P^{\prime}$ are two distinct occurring polygons containing $e$ whose interiors lie on the same side of $e$. Traverse both cycles starting from $e$ in the direction for which the interior stays, say, on the left. Let $x$ be the first vertex where the two traversals diverge. Let $y$ (resp. $y^{\prime}$) be the next vertex after $x$ along $P$ (resp. $P^{\prime}$), so $y\neq y^{\prime}$. By planarity and the fact that both interiors lie on the same side, one of $y,y^{\prime}$ must lie strictly inside the other polygon while being adjacent to $x$ on its boundary. Hence that vertex belongs to $\partial^{V,\circ}$ of the other polygon and therefore must be closed on the event that the other polygon occurs. But it also lies on an occurring polygon, so it must be open. This contradiction proves uniqueness for each choice of interior side and thus (7.3).
Taking expectations and using the lower bound on occurrence probabilities,
[…]
Therefore
$c_{n}(e)\leq 2\,[p(1-p)^{r}]^{-n},$
and taking $n$th roots and $\limsup_{n\to\infty}$ gives
$\gamma_{G}(e)\leq\frac{1}{p(1-p)^{r}}\qquad\text{for all }p\in(0,1).$
Optimizing over $p$ yields
$\gamma_{G}(e)\leq\min_{p\in[0,1]}\frac{1}{p(1-p)^{r}}=\frac{(1+r)^{1+r}}{r^{r}},$
as claimed. ∎
Lemma 7.4.
Let $G=(V,E)$ be an infinite, connected graph properly embedded into $\mathbb{R}^{2}$ with minimal degree at least $7$. Assume each finite face of $G$ is a triangle. Then
(7.4) $p_{c}^{site}(G_{*})\,\gamma_{G_{*}}(e)<1.$
Proof.
When each finite face of $G$ is a triangle, we have $G_{*}=G$. Applying Lemma 7.3 with $d=7$ gives $r=\frac{1}{2}$ and hence
(7.5) $\gamma_{G_{*}}(e)\leq\frac{(1+r)^{1+r}}{r^{r}}=\frac{(3/2)^{3/2}}{(1/2)^{1/2}}=\frac{3\sqrt{3}}{2}\approx 2.598.$
Let
$\alpha_{7}=\frac{1+\sqrt{5}}{2}.$
By [16, Theorem 2],
(7.6) $p_{c}^{site}(G_{*})=p_{c}^{site}(G)\leq\frac{2+\alpha_{7}}{4(1+\alpha_{7})}\approx 0.3455.$
Combining (7.5) and (7.6) yields (7.4). ∎
Remark 7.5.
Combining Lemmas 7.2 and 7.4, we may choose $\delta>0$ such that $p_{c}^{site}(G_{*})+\delta<1/\gamma_{G_{*}}(e)$ for all $e\in E_{*}$. In particular, Assumption 6.3 holds for graphs whose every finite face is a triangle and whose minimum degree is at least $7$.
8. Uniform Percolation with respect to Binary Trees
In this section we extend the near-$1$ non-uniqueness result from triangulations to general properly embedded planar graphs by (i) triangulating finite faces via edge additions and proving a stability statement under deleting the added edges, and (ii) invoking a binary-tree version of uniform percolation (in the sense of (8.1)) which avoids any bounded-degree hypothesis, in contrast to [27].
Throughout, $G=(V,E)$ is infinite, connected, properly embedded in $\mathbb{R}^{2}$, and has minimal vertex degree at least $7$. For each $x\in V$, Section 3 provides an embedded infinite rooted binary tree $T_{2,x}\subseteq G$.
Lemma 8.1.
Let $G=(V,E)$ be as above. For each $x\in V$, let $T_{2,x}$ be an infinite rooted binary tree embedded in $G$ with root $x$. Then for every $p\in(\tfrac{1}{2},1]$,
[…]
Proof.
From Section 3, we see that for any $x\in V$ we can find an infinite binary tree $T_{2,x}$ rooted at $x$ as a subgraph of $G$. Let $T_{2,x,N}\subset T_{2,x}$ be a depth-$N$ binary tree rooted at $x$.
Since the binary tree $T_{2}$ is self-similar and has $p_{c}^{site}(T_{2})=\frac{1}{2}$, we infer that for any $\epsilon>0$, there exists $N_{0}>0$, such that for any $N>N_{0}$
$\mathbb{P}_{p}(T_{2,x,N}\ \mathrm{intersects\ an\ infinite\ cluster\ in\ }T_{2,x})>1-\epsilon;$
in which the integer $N_{0}$ is independent of the root $x$ and the specific embedding of $T_{2,x}$ into $G$. Then the lemma follows since each infinite 1-cluster in $T_{2,x}$ must be a subset of an infinite 1-cluster on $G$. ∎
[…]
(8.2) $p_{c}^{\mathrm{site}}(\tilde{G})\leq p_{c}^{\mathrm{site}}(G).$
Exploration. Fix a total order on $V$. Given a site configuration $\eta\in\{0,1\}^{V}$ on $G$ and a vertex $x\in V$, we explore the $1$-cluster of $x$ in $G$ as follows. If $\eta(x)=0$, set $C_{0,G}(x)=\varnothing$ and $\partial C_{0,G}(x)=\varnothing$. If $\eta(x)=1$, set $C_{0,G}(x)=\{x\}$ and $\partial C_{0,G}(x)=\varnothing$. Inductively, at step $i\geq 1$: if there exists a vertex at $G$-distance 11 from $C_{i-1,G}(x)$ that is not in $\partial C_{i-1,G}(x)$, let $X_{i}$ be the first such vertex in the fixed order and define
$C_{i,G}(x)=\begin{cases}C_{i-1,G}(x)\cup\{X_{i}\},&\eta(X_{i})=1,\\ C_{i-1,G}(x),&\eta(X_{i})=0,\end{cases}\qquad\partial C_{i,G}(x)=\begin{cases}\partial C_{i-1,G}(x),&\eta(X_{i})=1,\\ \partial C_{i-1,G}(x)\cup\{X_{i}\},&\eta(X_{i})=0.\end{cases}$
If no such vertex exists, stop and set
$C_{G}(x):=C_{i-1,G}(x),\qquad\partial C_{G}(x):=\partial C_{i-1,G}(x).$
Then $C_{G}(x)=\bigcup_{i\geq 0}C_{i,G}(x)$ is exactly the $1$-cluster of $x$ in $G$ and $\partial C_{G}(x)=\bigcup_{i\geq 0}\partial C_{i,G}(x)$ is its (outer) vertex boundary. The same exploration applies to $\widetilde{G}$.
Lemma 8.2.
Let $G$ be as above and let $\tilde{G}$ be obtained by triangulating finite faces as above. For every $p\in(0,1]$ and every integer $N\geq 1$, $\mathbb{P}_{p}$-a.s. every infinite open cluster in $\tilde{G}$ contains a rooted binary tree of depth $N$ which uses only edges of $G$.
Proof.
The argument adapts the “finite-energy” multiscale proof of [27, Lemma 1.1], replacing balls by embedded binary trees (thus avoiding bounded-degree assumptions); see also the proof of [14, Theorem 2.3].
Fix $N\geq 3$ and $x\in V$. Let $C_{\tilde{G}}(x)$ denote the open cluster of $x$ in $\tilde{G}$. Let $E_{N,x}$ be the event that $C_{\tilde{G}}(x)$ is infinite but contains no rooted binary tree of depth $N$ in $G$. We show $\mathbb{P}_{p}(E_{N,x})=0$.
Write $B_{\tilde{G}}(x,r)$ for the ball of radius $r$ around $x$ in the graph metric of $\tilde{G}$. For $K\geq 1$, let $E_{N,x}^{K}$ be the event that $C_{\tilde{G}}(x)$ include a vertex in $\partial B_{\tilde{G}}(x,4NK)$ and contains no depth-$N$ rooted binary tree of $G$ in the ball $B_{\tilde{G}}(x,4NK)$.
Fix $K\geq 1$ and consider the outer annulus
$A:=B_{\tilde{G}}(x,4N(K+1))\setminus B_{\tilde{G}}(x,4NK+1).$
Let $\mathcal{T}_{N,K}$ be the set consisting of all the binary tree of depth $N$ with all the vertices in $A$. For each configuration in $E_{N,x}^{K}$ let $y$ be the minimal vertex in $C_{\tilde{G}}(x)\cap\partial B_{\tilde{G}}(x,4NK)$.
Do an exploration on the graph $G_{K}:=\tilde{G}\setminus B_{\tilde{G}}(x,4NK)$ starting from the minimal neighboring vertex of $y$ on $G_{K}$. Let $\tau_{K}$ be the first step in the exploration process such that the next unexplored vertex is in a tree $T\in\mathcal{T}_{N,K}$. It follows that
${\mathbb{P}}_{p}(C_{\tilde{G}}(x)\ \mathrm{contains\ a\ tree\ in\ }\mathcal{T}_{N,K}|E_{N,x}^{K}\cap\{\tau_{K}<\infty\})\geq p^{2^{N+1}-1}$
Note that
$E_{N,x}^{K+1}\subset[E_{N,x}^{K}\cap\{\tau_{K}<\infty\}]$
In particular, if $C_{\tilde{G}}(x)$ contains a vertex on $\partial B_{\tilde{G}}(x,4N(K+1))$, then it contains a vertex on $\partial B_{\tilde{G}}(x,4NK+2N)$, every neighbor of which is a root of a tree in $\mathcal{T}_{N,K}$, and therefore $\tau_{K}$ must be less than the first hitting time of $\partial B_{\tilde{G}}(x,4NK+2N)$; the latter is finite when $C_{\tilde{G}}(x)$ contains a vertex on $\partial B_{\tilde{G}}(x,4N(K+1))$.
By construction, at time $\tau_{K}$,the states of vertices in $V_{T}$ are not revealed yet; hence, conditional on the exploration history up to $\tau_{K}$, they remain i.i.d. Bernoulli($p$). Therefore the probability that all vertices of $T$ are open equals $p^{2^{N+1}-1}$. Then
(8.3) ${\mathbb{P}}_{p}(E_{N,x}^{K+1}\cap\{\tau_{K+1}<\infty\}|E_{N,x}^{K}\cap\{\tau_{K}<\infty\})$
$\leq 1-{\mathbb{P}}_{p}(C_{\tilde{G}}(x)\ \mathrm{contains\ a\ tree\ in\ }\mathcal{T}_{N,{K}}|E_{N,x}^{K}\cap\{\tau_{K}<\infty\})\leq 1-p^{2^{N+1}-1}$
Note that $E_{N,x}=\cap_{K=1}^{\infty}[E_{N,x}^{K}\cap\{\tau_{K}<\infty\}]$. The lemma follows from (8.3) since $p^{2^{N+1}-1}$ is independent of $K$, and the fact that
${\mathbb{P}}_{p}(E_{N,x})$ $=\lim_{k\rightarrow\infty}{\mathbb{P}}_{p}(E_{N,x}^{K}\cap\{\tau_{K}<\infty\})$
$\leq\lim_{K\rightarrow\infty}\prod_{i=2}^{K}{\mathbb{P}}_{p}(E_{N,x}^{i}\cap\{\tau_{i}<\infty\}|E_{N,x}^{i-1}\cap\{\tau_{i-1}<\infty\})$
$\leq\lim_{K\rightarrow\infty}(1-p^{2^{N+1}-1})^{K}=0.$
Lemma 8.3.
Let $G$ be as above and let $\tilde{G}$ be the triangulation of finite faces. Then there exists $\delta>0$ such that for every
[…]
$\mathbb{P}_{p}$-a.s. there are infinitely many infinite open clusters in $G$.
Proof.
If $G=\tilde{G}$ (i.e. every finite face is a triangle), the conclusion is exactly the near-$1$ non-uniqueness result proved for triangulations in the previous section.
Assume now $G\neq\tilde{G}$. For triangulations, there exists $\delta>0$ such that for every $p$ in the stated interval, $\mathbb{P}_{p}$-a.s. $\tilde{G}$ has infinitely many infinite open clusters. It therefore suffices to prove the following stability statement:
Claim. For such $p$, $\mathbb{P}_{p}$-a.s. every infinite open cluster of $\tilde{G}$ contains an infinite open cluster of $G$.
Fix $p$ in the stated interval. Note that $p>1/2$ since $p_{c}^{\mathrm{site}}(\tilde{G})<1/2$ (Section 3), so Lemma 8.1 applies.
Let $E_{x}$ be the event that $C_{\tilde{G}}(x)$ is infinite but contains no infinite open cluster of $G$, and let $E=\bigcup_{x\in V}E_{x}$. We show $\mathbb{P}_{p}(E)=0$. For a set $C\subseteq V$, define
[…]
For $x\in V$, split $E_{x}$ into $E_{x}^{f}:=E_{x}\cap\{W(C_{\tilde{G}}(x))<\infty\}$ and $E_{x}^{\infty}:=E_{x}\cap\{W(C_{\tilde{G}}(x))=\infty\}$, and set $E^{f}:=\bigcup_{x}E_{x}^{f}$, $E^{\infty}:=\bigcup_{x}E_{x}^{\infty}$. Then $E=E^{f}\cup E^{\infty}$.
Step 1: $\mathbb{P}_{p}(E^{0})=0$ where $E^{0}:=\bigcup_{x}(E_{x}\cap\{W=0\})$. Fix $x$. We use a duplication trick with two independent Bernoulli($p$) configurations $\omega^{\prime},\omega^{\prime\prime}$.
Explore the open cluster $C^{\prime}_{\tilde{G}}(x)$ in $\omega^{\prime}$ until either (i) the exploration finishes (so $C^{\prime}_{\tilde{G}}(x)$ is finite), or (ii) the explored open set contains a rooted binary tree of depth $N$ in $G$. In case (ii), let $Y$ be the root of the first such tree explored and denote this depth-$N$ tree by $T_{2,Y,N}$; let $S$ be the (finite) set of revealed vertices (explored open vertices together with their explored $\tilde{G}-$boundary).
Define a new configuration $\omega$ by setting $\omega=\omega^{\prime}$ on $S$ and $\omega=\omega^{\prime\prime}$ on $V\setminus S$. Then $\omega$ is again i.i.d. Bernoulli($p$).
Let $\widetilde{F}$ be the event that case (ii) occurs and that $T_{2,Y,N}$ intersects an infinite open cluster of $G$ in the independent configuration $\omega^{\prime\prime}$. On the event $E_{x}\cap\{W=0\}$ for $\omega$, the cluster $C_{\tilde{G}}(x)$ is infinite; hence the exploration cannot have terminated in case (i) (otherwise the revealed $\tilde{G}-$boundary would block any further growth), so case (ii) occurs. Moreover, if $\widetilde{F}$ occurs, then either $C_{\tilde{G}}(x)$ contains an infinite open $G$-cluster (if the intersection occurs through an open vertex of $C_{\tilde{G}}(x)$), or else some $\tilde{G}-$boundary vertex of $C_{\tilde{G}}(x)$ is $G$-adjacent to an infinite open $G$-cluster (because removing the finite set $S$ from an infinite $\omega^{\prime\prime}$-cluster leaves an infinite component adjacent to $S$). In either case, we cannot be in $E_{x}\cap\{W=0\}$. Thus
$\mathbb{P}_{p}(E_{x}\cap\{W=0\})\ \leq\ \mathbb{P}_{p}(\widetilde{F}^{\,c}\mid\text{case (ii)}).$
Conditioning on $Y$ and the choice of $T_{2,Y,N}$ and using independence of $\omega^{\prime}$ and $\omega^{\prime\prime}$, Lemma 8.1 implies that the right-hand side tends to 0 as $N\to\infty$. Hence $\mathbb{P}_{p}(E_{x}\cap\{W=0\})=0$ for all $x$, and therefore $\mathbb{P}_{p}(E^{0})=0$.
Step 2: $\mathbb{P}_{p}(E^{f})=0$. If $\mathbb{P}_{p}(E^{f})>0$, then for some fixed origin $v_{0}\in V$ there exists $M<\infty$ such that the event $E^{f,M}$ occurs with positive probability, where $E^{f,M}$ is the event that there exists an infinite open cluster $\xi$ in $\tilde{G}$ which contains no infinite open $G$-cluster and such that every vertex counted by $W(\xi)$ lies inside $B_{\tilde{G}}(v_{0},M)$.
Now couple two independent configurations $\omega^{\prime},\omega^{\prime\prime}$ and define $\eta$ by taking $\eta=\omega^{\prime}$ on $B_{\tilde{G}}(v_{0},2M)$ and $\eta=\omega^{\prime\prime}$ on its complement. On the event that $\omega^{\prime}\equiv 0$ on $B_{\tilde{G}}(v_{0},2M)$ and that $E^{f,M}$ occurs in $\omega^{\prime\prime}$, the configuration $\eta$ belongs to $E^{0}$ (all potential $W$-neighbors are killed inside the $2M$-ball). Hence $\mathbb{P}_{p}(E^{0})>0$, contradicting Step 1. Therefore $\mathbb{P}_{p}(E^{f})=0$.
Step 3: $\mathbb{P}_{p}(E^{\infty})=0$. Fix $x$. Assume we have explored $C_{G}(x)$ and $\partial C_{G}(x)$. If $|C_{G}(x)|=\infty$, then the infinite 1-cluster at $x$ in $\tilde{G}$ contains an infinite 1-cluster $C_{G}(x)$ in $G$.
Now assume $|C_{G}(x)|<\infty$. We explore $C_{\tilde{G}}(x)$ by growing it from $C_{G}(x)$. Set
$C_{0,\tilde{G}}(x):=C_{G}(x),\qquad\partial C_{0,\tilde{G}}(x)=\partial C_{G}(x),$
and iteratively reveal $\tilde{G}-$neighbors of $C_{i-1,\tilde{G}}(x)$ not in $\partial C_{i-1,\tilde{G}}(x)$ in a fixed order: if the next vertex $X_{i}$ is closed, put it into $\partial C_{i,\tilde{G}}(x)$; if it is open, adjoin it to the explored set. This yields $C_{\tilde{G}}(x)=\bigcup_{i}C_{i,\tilde{G}}(x)$.
Let $I_{1}<I_{2}<\cdots$ be the indices $i$ such that $X_{i}\notin[C_{G}(x)\cup\partial C_{G}(x)]$ and $X_{i}$ is $G$-adjacent to an infinite open $G$-cluster. By independence of vertex states, the random variables $\omega(X_{I_{1}}),\omega(X_{I_{2}}),\dots$ are i.i.d. Bernoulli($p$). On the event $E_{x}^{\infty}$, there are infinitely many such indices and necessarily $\omega(X_{I_{j}})=0$ for all $j$ (otherwise an infinite open $G$-cluster would be absorbed into $C_{\tilde{G}}(x)$), which has probability $\lim_{n\to\infty}(1-p)^{n}=0$. Hence $\mathbb{P}_{p}(E_{x}^{\infty})=0$ for each $x$, and by countable union $\mathbb{P}_{p}(E^{\infty})=0$.
Combining Steps 1–3 gives $\mathbb{P}_{p}(E)=0$, proving the Claim and hence the lemma. ∎
Lemma 8.4.
Let $G=(V,E)$ be as above and couple percolation by i.i.d. $U(v)\sim\mathrm{Unif}[0,1]$. For $p\in[0,1]$, let $V_{p}=\{v:U(v)\leq p\}$ and $G_{p}$ be the induced subgraph. Assume $\tfrac{1}{2}<p_{1}<p_{2}\leq 1$ and that $G$ has uniform percolation at level $p_{1}$ in the sense of Lemma 8.1. Then almost surely every infinite cluster of $G_{p_{2}}$ contains at least one infinite cluster of $G_{p_{1}}$.
Proof.
This is proved by the same duplication/exploration argument as [27, Theorem 1.1], replacing “balls” by “depth-$N$ embedded binary trees” and invoking Lemma 8.1 instead of ball-uniformity. The bounded-degree assumption in [27] is only used to work with balls; the binary-tree version avoids this. We omit the repeated details. ∎
[…]
Acknowledgements. ZL thanks Russell Lyons for comments. ZL acknowledges support from the National Science Foundation DMS 1608896 and Simons Foundation grant 638143.